### 27 Feb 2024 v2

# Non-projective two-weight codes

Abstract.
It has been known since the 1970’s that the difference of the non-zero weights of a projective $\mathbb{F}_{q}$-linear two-weight has to be a power of the characteristic of the underlying field. Here we study non-projective two-weight codes and e.g. show the same result under mild extra conditions. For small dimensions we give exhaustive enumerations of the feasible parameters in the binary case.
Keywords: Linear codes, two-weight codes, two-character sets
MSC: 94B05; 05B25, 68R01
1. Introduction
It has been known since the 1970’s that the two non-zero weights of a projective $\mathbb{F}_{q}$-linear two-weight code $C$ can be written as $w_{1}=up^{t}$ and $w_{2}=(u+1)p^{t}$, where $u\in\mathbb{N}_{\geq 1}$ and $p$ is the characteristic of the underlying finite field $\mathbb{F}_{q}$, see [19, Corollary 2]. So, especially the weight difference $w_{2}-w_{1}$ is a power of the characteristic $p$. Here we want to consider $\mathbb{F}_{q}$-linear two-weight codes $C$ with non-zero weights $w_{1}<w_{2}$ which are not necessarily projective. In [12] it was observed that if $w_{2}-w_{1}$ is not a power of the characteristic $p$, then the code $C$ has to be non-projective, which settles a question in [36]. Here we prove the stronger statement that $C$ is repetitive, i.e., $C$ is the $l$-fold repetition of a smaller two-weight code $C^{\prime}$, where $l$ is the largest factor of $w_{2}-w_{1}$ that is coprime to the field size $q$, if $C$ does not have full support, c.f. [11]. Moreover, if a two-weight code $C$ is non-repetitive and does not have full support, then its two non-zero weights can be written as $w_{1}=up^{t}$ and $w_{2}=(u+1)p^{t}$, where again $p$ is the characteristic of the underlying finite field $\mathbb{F}_{q}$, see Theorem 5.14
Constructions for projective two-weight codes can be found in the classical survey paper [16]. Many research papers considered these objects since they e.g. yield strongly regular graphs and we refer to [14] for a corresponding monograph on srgs. For a few more recent papers on constructions for projective two-weight codes we refer e.g. to [28, 32, 37, 41]. In e.g. [37] the author uses geometric language and speaks of constructions for two-character sets, i.e., sets of points in a projective space $\operatorname{PG}(k-1,q)$ with just two different hyperplane multiplicities, call them $s$ and $t$. In general each (full-length) linear code is in one-to-one correspondence to a (spanning) multiset of points in some projective space $\operatorname{PG}(k-1,q)$. Here we will also mainly use the geometric language and consider a few general constructions for two-character multisets of points corresponding to two-weight codes (possibly non-projective). For each subset $\overline{\mathcal{H}}$ of hyperplanes in $\operatorname{PG}(k-1,q)$ we construct a multiset of points $\mathcal{M}(\overline{\mathcal{H}})$ such that all hyperplanes $H\in\overline{\mathcal{H}}$ have the same multiplicity, say $s$, and also all other hyperplanes $H\notin\overline{\mathcal{H}}$ have the same multiplicity, say $t$. Actually, we characterize the full set of such multisets with at most two different hyperplane multiplicities given $\overline{\mathcal{H}}$, see Theorem 5.11 and Theorem 5.13. Using this correspondence we have classified all two-weight codes up to symmetry for small parameters. For projective two-weight codes such enumerations can be found in [7].
Brouwer and van Eupen give a correspondence between arbitrary projective codes and arbitrary two-weight codes via the so-called BvE-dual transform. The correspondence can be said to be 1-1, even though there are choices for the involved parameters to be made in both directions. In [13] the dual transform was e.g. applied to the unique projective $[16,5,9]_{3}$-code. For parameters $\alpha=\tfrac{1}{3}$, $\beta=-3$ the result is a $[69,5,45]_{3}$ two-weight code and for $\alpha=-\tfrac{1}{3}$, $\beta=5$ the results is a $[173,5,108]_{3}$ two-weight code. This resembles the fact that we have some freedom when constructing a two-weight code from a given projective code, e.g. we can take complements or add simplex codes of the same dimension. Our obtained results may be rephrased in the language of the BvE-dual transform by restricting to a canonical choice of the involved parameters. For further literature on the dual transform see e.g. [5, 9, 13, 38]. For a variant that is rather close to our presentation we refer to [10].
With respect to further related literature we remark that a special subclass of (non-projective) two-weight codes was completely characterized in [30]. A conjecture by Vega [39] states that all two-weight cyclic codes are the “known” ones, c.f. [22]. Another stream of literature considers the problem whether all projective two-weight codes that have the parameters of partial $k$-spreads indeed have to be partial $k$-spreads. Those results can be found in papers considering extendability results for partial $k$-spreads or classifying minihypers, see e.g. [24]. Several non-projective two-weight codes appear also as minimum length examples for divible minimal codes [35]. [^1] Minimal codes are linear codes where all non-zero codewords are minimal, i.e., whose support is not properly contained in the support of another codeword. Two-weight codes have also been considered over rings instead of finite fields, see e.g. [15].
The remaining part of this paper is structured as follows. In Section 2 we introduce the necessary preliminaries for linear two-weight codes and their geometric counterpart called two-character multisets in projective spaces. In general multisets of points, corresponding to general linear codes, can be described via so-called characteristic functions and we collect some of their properties in Section 3. Examples and constructions for two-character multisets are given in Section 4. In Section 5 we present our main results. We close with enumeration results for two-character multisets in $\operatorname{PG}(k-1,q)$ for small parameters in Section 6. We will mainly use geometric language and arguments. For the for the ease of the reader we only use elementary arguments and give (almost) all proof details. In an appendix we list all candidates for sets of parameters of two-character sets in $\operatorname{PG}(k-1,2)$ for $4\leq k\leq 12$ and state corresponding constructions from the literature. A few cases are left open for $k\in\{11,12\}$.
2. Preliminaries
[…]
Now let $C$ be a full-length $[n,k]_{q}$-code with generator matrix $G$. Each column $g$ of $G$ is an element of $\mathbb{F}_{q}^{k}$ and since $g\neq\mathbf{0}$ we can consider $\langle g\rangle$ as point in the projective space $\operatorname{PG}(k-1,q)$. Using the geometric language we call $1$-, $2$-, $3$-, and $(k-1)$-dimensional subspaces of $\mathbb{F}_{q}^{k}$ points, lines, planes, and hyperplanes in $\operatorname{PG}(k-1,q)$. Instead of an $l$-dimensional space we also speak of an $l$-space. By $\mathcal{P}$ we denote the set of points and by $\mathcal{H}$ we denote the set of hyperplanes of $\operatorname{PG}(k-1,q)$ whenever $k$ and $q$ are clear from the context. A multiset of points in $\operatorname{PG}(k-1,q)$ is a mapping $\mathcal{M}\colon\mathcal{P}\to\mathbb{N}$, i.e., to each point $P\in\mathcal{P}$ we assign its multiplicity $\mathcal{M}(P)\in\mathbb{N}$. By #$\#\mathcal{M}=\sum_{P\in\mathcal{P}}\mathcal{M}(P)$ we denote the cardinality of $\mathcal{M}$. The support $\operatorname{supp}(\mathcal{M})$ is the set of all points with non-zero multiplicity. We say that $\mathcal{M}$ is spanning if the set of points in the support of $\mathcal{M}$ span $\operatorname{PG}(k-1,q)$. Clearly permuting columns of a generator matrix $G$ or multiplying some columns with non-zero elements in $\mathbb{F}_{q}^{\star}:=\mathbb{F}_{q}\backslash\{0\}$ yields an equivalent code. Besides that we get an one-to-one correspondence between full length $[n,k]_{q}$-codes and spanning multisets of points $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$ with cardinality #$\#\mathcal{M}=n$. Moreover, two linear $[n,k]_{q}$-codes $C$ and $C^{\prime}$ are equivalent iff their corresponding multisets of points $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are. For details we refer e.g. to [21]. A linear code $C$ is projective iff its corresponding multiset of points satisfies $\mathcal{M}(P)\in\{0,1\}$ for all $P\in\mathcal{P}$. We also speak of a set of points in this case. The multisets of points with $\mathcal{M}(P)=0$ for all $P\in\mathcal{P}$ are called trivial.
Geometrically, for a non-zero codeword $c\in C$ the set $c\cdot\mathbb{F}_{q}^{\star}$ corresponds to a hyperplane $H\in\mathcal{H}$ and $\operatorname{wt}(c)=\#\mathcal{M}-\mathcal{M}(H)$, where we extend the function $\mathcal{M}$ additively. i.e., $\mathcal{M}(S):=\sum_{P\in S}\mathcal{M}(P)$ for every subset $S\subseteq\mathcal{P}$ of points. We call $\mathcal{M}(H)$ the multiplicity of hyperplane $H\in\mathcal{H}$ and have $\mathcal{M}(V)=\#\mathcal{M}$ for the entire ambient space $V:=\mathcal{P}$. The number of hyperplanes $\#\mathcal{H}$, as well as the number of points #$\#\mathcal{P}$, in $\operatorname{PG}(k-1,q)$ is given by $[k]_{q}:=\tfrac{q^{k}-1}{q-1}$. A two-character multiset is a multiset of points $\mathcal{M}$ such that exactly two different hyperplane multiplicities $\mathcal{M}(H)$ occur. I.e., a multiset of points $\mathcal{M}$ is a two-character multiset iff its corresponding code $C$ is a two-weight code. If $\mathcal{M}$ actually is a set of points, i.e. if we have $\mathcal{M}(P)\in\{0,1\}$ for all points $P\in\mathcal{P}$, then we speak of a two-character set. We say that an $[n,k]_{q}$-code $C$ is $\Delta$-divisible if the weights of all codewords are divisible by $\Delta$. A multiset of points $\mathcal{M}$ is called $\Delta$-divisible if the corresponding linear code is. More directly, a multiset of points $\mathcal{M}$ is $\Delta$-divisible if we have $\mathcal{M}(H)\equiv\#\mathcal{M}\pmod{\Delta}$ for all $H\in\mathcal{H}$.
A one-weight code is an $[n,k]_{q}$-code $C$ such that all non-zero codewords have the same weight $w$. One-weight codes have been completely classified in [3] and are given by repetitions of so-called simplex codes. Geometrically, the multiset of points $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$ corresponding to a one-weight code $C$ satisfies $\mathcal{M}(P)=l$ for all $P\in\mathcal{P}$. I.e., we have #$\#\mathcal{M}=n=[k]_{q}\cdot l$, $\mathcal{M}(H)=[k-1]_{q}\cdot l$ for all $H\in\mathcal{H}$, and $w=\#\mathcal{M}-\mathcal{M}(H)=q^{k-1}\cdot l$. We say that a linear $[n,k]_{q}$-code $C$ is repetitive if it is the $l$-fold repetition of an $[n/l,k]_{q}$-code $C^{\prime}$, where $l>1$, and non-repetitive otherwise. A given multiset of points $\mathcal{M}$ is called repeated if its corresponding code is. More directly, a non-trivial multiset of points $\mathcal{M}$ is repeated iff the greatest common divisor of all point multiplicities is larger than one. We say that a multiset of points $\mathcal{M}$ or its corresponding linear code $C$ has full support iff $\operatorname{supp}(\mathcal{M})=\mathcal{P}$, i.e., if $\mathcal{M}(P)>0$ for all $P\in\mathcal{P}$. So, for each non-repetitive one-weight code $C$ with length $n$, dimension $k$, and non-zero weight $w$ we have $n=[k]_{q}$ and $w=q^{k-1}$. Each non-trivial one-weight code, i.e., one with dimension at least $1$, has full support. The aim of this paper is to characterize the possible parameters of non-repetitive two-weight codes (with or without full support). For the correspondence between $[n,k]_{q}$-codes and multisets of points $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$ we have assumed that $\mathcal{M}$ is spanning. If $\mathcal{M}$ is not spanning, then there exists a hyperplane containing the entire support, so that $\mathcal{M}$ is two-character multiset iff $\mathcal{M}$ induces a one-character multiset in the span of $\operatorname{supp}(\mathcal{M})$ cf. Proposition 4.1. The structure of the set of all two-character multisets where the larger hyperplane multiplicity is attained for a prescribed subset of the hyperplanes is considered in Section 5.
3. Characteristic functions
Fixing the field size $q$ and the dimension $k$ of the ambient space, a multiset of points in $\operatorname{PG}(k-1,q)$ is a mapping $\mathcal{M}\colon\mathcal{P}\to\mathbb{N}$. By $\mathcal{F}$ we denote the $\mathbb{Q}$-vector space consisting of all functions $F\colon\mathcal{P}\to\mathbb{Q}$, where addition and scalar multiplication is defined pointwise. I.e., $(F_{1}+F_{2})(P):=F_{1}(P)+F_{2}(P)$ and $(x\cdot F_{1})(P):=x\cdot F_{1}(P)$ for all $P\in\mathcal{P}$, where $F_{1},F_{2}\in\mathcal{F}$ and $x\in\mathbb{Q}$ are arbitrary. For each non-empty subset $S\subseteq\mathcal{P}$ the characteristic function $\chi_{S}$ is defined by $\chi_{S}(P)=1$ if $P\in S$ and $\chi_{S}(P)$ otherwise. Clearly the set of functions $\chi_{P}$ for all $P\in\mathcal{P}$ forms a basis of $\mathcal{F}$ for ambient space $\operatorname{PG}(k-1,q)$ for all $k\geq 1$. Note that there are no hyperplanes if $k=1$ and hyperplanes coincide with points for $k=2$. We also extend the functions $F\in\mathcal{F}$ additively, i.e., we set $F(S)=\sum_{P\in S}F(P)$ for all $S\subseteq\mathcal{P}$. Our next aim is to show the well-known fact that also the set of functions $\chi_{H}$ for all hyperplanes $H\in\mathcal{H}$ forms a basis of $\mathcal{F}$ for ambient space $\operatorname{PG}(k-1,q)$ for all $k\geq 2$. In other words, also $\mathcal{M}(P)$ can be reconstructed from the $\mathcal{M}(H)$:
Lemma 3.1.
Let $\mathcal{M}\in\mathcal{F}$ for ambient space $\operatorname{PG}(k-1,q)$, where $k\geq 2$. Then, we have
$\mathcal{M}(P)=\sum_{H\in\mathcal{H}\,:\,P\in H}\frac{1}{[k-1]_{q}}\cdot\mathcal{M}(H)+\sum_{H\in\mathcal{H}\,:\,P\notin H}\frac{1}{q^{k-1}}\cdot\left(\frac{1}{[k-1]_{q}}-1\right)\cdot\mathcal{M}(H)$ (1)
for all points $P\in\mathcal{P}$.
  • Proof. W.l.o.g. we assume $k\geq 3$. Since each point $P^{\prime}\in\mathcal{P}$ is contained in $[k-1]_{q}$ of the #$\#\mathcal{H}=[k]_{q}$ hyperplanes and each point $P^{\prime}\neq P$ is contained in $[k-2]_{q}$ of the $[k-1]_{q}$ hyperplanes that contain $P$, we have
    $\sum_{H\in\mathcal{H}\,:\,P\in H}\mathcal{M}(H)=[k-2]_{q}\cdot|\mathcal{M}|+\left([k-1]_{q}-[k-2]_{q}\right)\mathcal{M}(P)=[k-2]_{q}\cdot|\mathcal{M}|+q^{k-2}\mathcal{M}(P)$
    so that
    $\sum_{H\in\mathcal{H}\,:\,P\in H}\mathcal{M}(H)-\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\sum_{H\in\mathcal{H}}\mathcal{M}(H)=q^{k-2}\mathcal{M}(P)$
    using $[k-1]_{q}\cdot\#\mathcal{M}=\sum_{H\in\mathcal{H}}\mathcal{M}(H)$. Thus, we can conclude the stated formula using
    $\frac{1}{q^{k-2}}\cdot\left(1-\frac{[k-2]_{q}}{[k-1]_{q}}\right)=\frac{1}{q^{k-2}}\cdot\frac{[k-1]_{q}-[k-2]_{q}}{[k-1]_{q}}=\frac{1}{[k-1]_{q}}$
    and
    $-\frac{[k-2]_{q}}{[k-1]_{q}\cdot q^{k-2}}=\frac{1-[k-1]_{q}}{[k-1]_{q}\cdot q^{k-1}}=\frac{1}{q^{k-1}}\cdot\left(\frac{1}{[k-1]_{q}}-1\right).$
    $\square$
As an example we state that in $\operatorname{PG}(3-1,2)$ we have
$\mathcal{M}(P)=\frac{1}{3}\cdot\sum_{H\in\mathcal{H}\,:\,P\in H}\mathcal{M}(H)-\frac{1}{6}\cdot\sum_{H\in\mathcal{H}\,:\,P\notin H}\mathcal{M}(H).$
Lemma 3.2.
Let $\mathcal{M}\in\mathcal{F}$ for ambient space $\operatorname{PG}(k-1,q)$, where $k\geq 2$. Then there exist $\alpha_{H}\in\mathbb{Q}$ for all hyperplanes $H\in\mathcal{H}$ such that
$\mathcal{M}=\sum_{H\in\mathcal{H}}\alpha_{H}\cdot\chi_{H}.$ (2)
Moreover, the coefficients $\alpha_{H}$ are uniquely determined by $\mathcal{M}$.
  • Proof. From
    $\sum_{H\in\mathcal{H}\,:\,P\in H}\chi_{H}-\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\sum_{H\in\mathcal{H}}\chi_{H}=q^{k-2}\cdot\chi_{P}$
    for each point $P\in\mathcal{P}$ and
    $\mathcal{M}=\sum_{P\in\mathcal{P}}\mathcal{M}(P)\cdot\chi_{P}$
    we conclude the existence of the $\alpha_{H}\in\mathbb{Q}$. Since the functions $\left(\chi_{P}\right)_{P\in\mathcal{P}}$ form a basis of the $\mathbb{Q}$-vector space $\mathcal{F}$ , which is also generated by the functions $\left(\chi_{H}\right)_{H\in\mathcal{H}}$, counting #$\#\mathcal{P}=[k]_{q}=\#\mathcal{H}$ yields that also $\left(\chi_{H}\right)_{H\in\mathcal{H}}$ forms a basis and the coefficients $\alpha_{H}$ are uniquely determined by $\mathcal{M}$. $\square$
If $\mathcal{M}\in\mathcal{F}$ is given by the representation
$\mathcal{M}=\sum_{P\in\mathcal{P}}\alpha_{P}\cdot\chi_{P}$
with $\alpha_{P}\in\mathbb{Q}$ we can easily decide whether $\mathcal{M}$ is a multiset of points. The necessary and sufficient conditions are given by $\alpha_{P}\in\mathbb{N}$ for all $P\in\mathcal{P}$ (including the case of a trivial multiset of points). If a multiset of points is characterized by coefficients $\alpha_{H}$ for all hyperplanes $H\in\mathcal{H}$ as in Lemma 3.2 then some $\alpha_{H}$ may be fractional or negative. For two-character multisets we will construct a different unique representation involving the characteristic functions $\chi_{H}$ of hyperplanes, see Theorem 5.11.
Let us state a few observations about operations for multisets of points that yield multisets of points again.
Lemma 3.3.
For two multisets of points $\mathcal{M}_{1},\mathcal{M}_{2}$ of $\operatorname{PG}(k-1,q)$ and each non-negative integer $n\in\mathbb{N}$ the functions $\mathcal{M}_{1}+\mathcal{M}_{2}$ and $n\cdot\mathcal{M}_{1}$ are multisets of points of $\operatorname{PG}(k-1,q)$.
In order to say something about the subtraction of multisets of points we denote the minimum point multiplicity of a multiset of points $\mathcal{M}$ by $\mu(\mathcal{M})$ and the maximum point multiplicity by $\gamma(\mathcal{M})$. Whenever $\mathcal{M}$ is clear from the context we also just write $\mu$ and $\gamma$ instead of $\mu(\mathcal{M})$ and $\mu(\gamma)$.
Lemma 3.4.
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two multisets of points of $\operatorname{PG}(k-1,q)$. If $\mu(\mathcal{M}_{1})\geq\gamma(\mathcal{M}_{2})$, then $\mathcal{M}_{1}-\mathcal{M}_{2}$ is a multiset of points of $\operatorname{PG}(k-1,q)$.
Definition 3.5.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$. If $l$ is an integer with $l\geq\gamma(\mathcal{M})$, then the $l$-complement $\mathcal{M}^{l-C}$ of $\mathcal{M}$ is defined by $\mathcal{M}^{l-C}(P):=l-\mathcal{M}(P)$ for all points $P\in\mathcal{P}$.
One can easily check that $\mathcal{M}^{l-C}$ is a multiset of points with cardinality $l\cdot[k]_{q}-\#\mathcal{M}$, maximum point multiplicity $\gamma\!\left(\mathcal{M}^{l-C}\right)=l-\mu(\mathcal{M})$, and minimum point multiplicity $\mu\!\left(\mathcal{M}^{l-C}\right)=l-\gamma(\mathcal{M})$. Using characteristic functions we can write $\mathcal{M}^{l-C}=l\cdot\chi_{V}-\mathcal{M}$, where $V=\mathcal{P}$ denotes the set of all points of the ambient space.
Given an arbitrary function $\mathcal{M}\in\mathcal{F}$ there always exist $\0\alpha\in\mathbb{Q}\backslash\{0\}$ and $\beta\in\mathbb{Z}$ such that $\alpha\cdot\mathcal{M}+\beta\cdot\chi_{V}$ is a multiset of points.
4. Examples and constructions for two-character multisets
The aim of this section is to list a few easy constructions for two-character multisets of points $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$. We will always abbreviate $n=\#\mathcal{M}$ and denote the two occurring hyperplane multiplicities by $s$ and $t$, where we assume $s>t$ by convention.
Proposition 4.1.
For integers $1\leq l<k$ let $L$ be an arbitrary $l$-space in $\operatorname{PG}(k-1,q)$. Then $\chi_{L}$ is a two-character set with $n=[l]_{q}$, $\gamma=1$, $\mu=0$, $s=[l]_{q}$, and $t=[l-1]_{q}$.
Note that for the case $l=k$ we have the one-character set $\chi_{V}$, which can be combined with any two-character multiset.
Lemma 4.2.
Let $\mathcal{M}$ be a two-character multiset of points in $\operatorname{PG}(k-1,q)$. Then, for each integer $0\leq a\leq\mu(\mathcal{M})$, each $b\in N$, and each integer $c\geq\gamma(\mathcal{M})$ the functions $\mathcal{M}-a\cdot\chi_{V}$, $\mathcal{M}+b\cdot\chi_{V}$, $b\cdot\mathcal{M}$, and $c\cdot\chi_{V}-\mathcal{M}$ are two-character multisets of points.
For the first and the fourth construction we also spell out the implications for the parameters of a given two-character multiset:
Lemma 4.3.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}(H)\in\{s,t\}$ for every hyperplane $H\in\mathcal{H}$. If $\mathcal{M}(P)\geq l$ for every point $P\in\mathcal{P}$, i.e., $l\leq\mu(\mathcal{M})$, then $\mathcal{M}^{\prime}:=\mathcal{M}-l\cdot\chi_{V}$ is a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}^{\prime}(H)\in\{s-[k-1]_{q}\cdot l,t-[k-1]_{q}\cdot l\}$ for every hyperplane $H\in\mathcal{H}$.
Lemma 4.4.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}(H)\in\{s,t\}$ for every hyperplane $H\in\mathcal{H}$. If $\mathcal{M}(P)\leq u$, i.e. $\leq\gamma(\mathcal{M})$ for every point $P\in\mathcal{P}$, then the $u$-complement $\mathcal{M}^{\prime}:=u\cdot\chi_{V}-\mathcal{M}$ of $\mathcal{M}$ is a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}^{\prime}(H)\in\left\{u[k-1]-s,u[k-1]-t\right\}$ for every hyperplane $H\in\mathcal{H}$.
We can also use two (almost) arbitrary subspaces to construct two-character multisets:
Proposition 4.5.
Let $a\geq b\geq 1$ and $10\leq i\leq b-$ 1 be arbitrary integers, $A$ be an $a$-space and $B$ be an $b$-space with $\dim(A\cap B)=i$ in $\operatorname{PG}(k-1,q)$, where $k=a+b-i$, Then, $\mathcal{M}=\chi_{A}+q^{a-b}\cdot\chi_{B}$ satisfies $\mathcal{M}(H)\in\{s,t\}$ for all $H\in\mathcal{H}$, where $s=[a-1]_{q}+q^{a-b}\cdot[b-1]_{q}$ and $t=s+q^{a-1}$. If $i=0$, then $\gamma=q^{a-b}$ and $\gamma=q^{a-b}+1$ otherwise. In general, we have $n=[a]_{q}+q^{a-b}\cdot[b]_{q}$ and $\mu=0$.
  • Proof. For each $H\in\mathcal{H}$ we have $\mathcal{M}(H\cap A)\in\left\{[a-1]_{q},[a]_{q}\right\}$ and $\mathcal{M}(H\cap B)\in\left\{[b-1]_{q},[b]_{q}\right\}$. Noting that we cannot have both $\mathcal{M}(H\cap A)=[a]_{q}$ and $\mathcal{M}(H\cap B)=[b]_{q}$ we conclude $\mathcal{M}(H)\in\left\{[a-1]_{q}+q^{a-b}\cdot[b-1]_{q},[a]_{q}+q^{a-b}\cdot[b-1]_{q},[a-1]_{q}+q^{a-b}\cdot[b]_{q}\right\}=\{s,t\}$. $\square$
A partial $k$-spread is a set of $k$-spaces in $\operatorname{PG}(v-1,q)$ with pairwise trivial intersection.
Proposition 4.6.
Let $\mathcal{S}_{1},\dots,\mathcal{S}_{r}$ be a partial parallelism of $\operatorname{PG}(2k-1,q)$, i.e., the $\mathcal{S}_{i}$ are partial $k$-spreads that are pairwise disjoint. Then
$\mathcal{M}=\sum_{i=1}^{r}\sum_{S\in\mathcal{S}_{i}}\chi_{S}$
is a two-character multiset of $\operatorname{PG}(2k-1,q)$ with $n=r\cdot[k]_{q}$ and hyperplane multiplicities $s=r\cdot[k-1]_{q}$, $t=r\cdot[k-1]_{q}+q^{k-1}$, where $r=\sum_{i=1}^{r}\left|\mathcal{S}_{i}\right|$.
C.f. Example SU2 in [16]. Field changes work similarly as explained in [16, Section 6] for two-character sets.
Based on hyperplanes we can construct large families of two-character multisets:
Lemma 4.7.
Let $\emptyset\neq\mathcal{H}^{\prime}\subsetneq\mathcal{H}$ be a subset of the hyperplanes of $\operatorname{PG}(k-1,q)$, where $k\geq 3$, then
$\mathcal{M}=\sum_{H\in\mathcal{H}^{\prime}}\chi_{H}$ (3)
is a two-character multiset with $n=r[k-1]_{q}$, $s=r[k-2]_{q}+q^{k-2}$, and $t=r[k-2]_{q}$, where $r=\#\mathcal{H}^{\prime}$.
By allowing $\mathcal{H}^{\prime}$ to be a multiset of hyperplanes, we end up with $(\tau+1)$-character sets, where $\tau$ is the maximum number of occurrences of a hyperplane in $\mathcal{H}^{\prime}$.
Applying Lemma 4.3 yields:
Lemma 4.8.
Let $\emptyset\neq\mathcal{H}^{\prime}\subsetneq\mathcal{H}$ be a subset of the hyperplanes of $\operatorname{PG}(k-1,q)$, where $k\geq 3$. If each point $P\in\mathcal{P}$ is contained in at least $\mu\in N$ elements of $\mathcal{H}^{\prime}$, then
$\mathcal{M}=\sum_{H\in\mathcal{H}^{\prime}}\chi_{H}\,-\,\mu\cdot\chi_{V}$ (4)
is a two-character multiset with $n=r[k-1]_{q}-\mu[k]_{q}$, $s=r[k-2]_{q}+q^{k-2}-\mu[k-1]_{q}$ and $t=r[k-2]_{q}-\mu[k-1]_{q}$, where $r=|\mathcal{H}^{\prime}|$.
In some cases we obtain two-character multisets where all point multiplicities have a common factor $g>1$. Here we can apply the following general construction:
Lemma 4.9.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}(H)\in\{s,t\}$ for every hyperplane $H\in\mathcal{H}$. If $\mathcal{M}(P)\equiv 0\pmod{g}$ for every point $P\in\mathcal{P}$, then $\mathcal{M}^{\prime}:=\tfrac{1}{g}\cdot\mathcal{M}$ is a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}^{\prime}(H)\in\left\{\tfrac{1}{g}\cdot s,\tfrac{1}{g}\cdot t\right\}$ for every hyperplane $H\in\mathcal{H}$. Moreover, we have #$\#\mathcal{M}^{\prime}=\tfrac{1}{g}\cdot\#\mathcal{M}$, $\mu(\mathcal{M}^{\prime})=\tfrac{1}{g}\cdot\mu(\mathcal{M})$, and $\gamma(\mathcal{M}^{\prime})=\tfrac{1}{g}\cdot\gamma(\mathcal{M})$.
Interestingly enough, it will turn out that we can construct all two-character multisets by combining Lemma 4.7 with Lemma 4.2 and Lemma 4.9, see Theorem 5.11 and Theorem 5.13.
5. Geometric duals and sets of feasible parameters for two-character multisets
To each two-character multiset $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$, i.e., $\left\{\mathcal{M}(H)\,:\,H\in\mathcal{H}\right\}=\{s,t\}$ for some $s,t\in\mathbb{N}$ we can assign a set of points $\overline{\mathcal{M}}$ by using the geometric dual, i.e., interchanging hyperplanes and points. More precisely, fix a non-degenerated billinear form $\perp$ and consider pairs of points and hyperplanes $(P,H)$ that are perpendicular w.r.t. to $\perp$. We write $H=P^{\perp}$ for the geometric dual of a point. We define $\overline{\mathcal{M}}$ via ¯$\overline{\mathcal{M}}(P)=1$ iff $\mathcal{M}(H)=s$, where $H=P^{\perp}$, and ¯$\overline{\mathcal{M}}(P)=0$ otherwise, i.e., if $\mathcal{M}(H)=t$ [^2] A generalization of the notion of the geometric dual has been introduced by Brouwer and van Eupen [13] for linear codes and formulated for multisets of points by Dodunekov and Simonis [21]. Of course we have some freedom how we order $s$ and $t$. So, we may also write ¯$\overline{\mathcal{M}}(P)=\left(\mathcal{M}(H)-t\right)/(s-t)\in\{0,1\}$ for all $P\in\mathcal{P}$, where $H=P^{\perp}$. Noting the asymmetry in $s$ and $t$ we may also interchange the role of $s$ and $t$ or replace $\overline{\mathcal{M}}$ by its complement. Note that in principle several multisets of points with two hyperplane multiplicities can have the same corresponding point set $\overline{\mathcal{M}}$.
For the other direction we can start with an arbitrary set of points $\overline{\mathcal{M}}$, i.e., ¯$\overline{\mathcal{M}}(P)\in\{0,1\}$ for all $P\in\mathcal{P}$. The multiset of points with two hyperplane multiplicities $\mathcal{M}$ is then defined via $\mathcal{M}(H)=s$ if ¯$\overline{\mathcal{M}}(P)=1$, where $H=P^{\perp}$, and $\mathcal{M}(H)=t$ if ¯$\overline{\mathcal{M}}(P)=0$. I.e., we may set
$\mathcal{M}(H)=t+(s-t)\cdot\overline{\mathcal{M}}(H^{\perp}).$ (5)
While we have $\mathcal{M}(H)\in\mathbb{N}$ for all $s,t\in\mathbb{N}$, the point multiplicities $\mathcal{M}(P)$ induced by the hyperplane multiplicities $\mathcal{M}(H)$, see Lemma 3.1, are not integral or non-negative in general. For suitable choices of $s$ and $t$ they are, for others they are not.
Definition 5.1.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$. By $\mathbb{L}(\overline{\mathcal{M}})\subseteq\mathbb{N}^{2}$ we denote the set of all pairs $(s,t)\in\mathbb{N}^{2}$ with $s\geq t$ such that a multiset of points $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$ exists with $\mathcal{M}(H)=s$ if ¯$\overline{\mathcal{M}}(H^{\perp})=1$ and $\mathcal{M}(H)=t$ if ¯$\overline{\mathcal{M}}(H^{\perp})=0$ for all hyperplanes $H\in\mathcal{H}$.
Directly from Lemma 4.2 we can conclude:
Lemma 5.2.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$. If $(s,t)\in\mathbb{L}(\overline{\mathcal{M}})$, then we have
$\langle(s,t)\rangle_{\mathbb{N}}+\left\langle\left([k-1]_{q},[k-1]_{q}\right)\right\rangle_{\mathbb{N}}=\left\{\left(us+v[k-1]_{q},ut+v[k-1]_{q}\right)\,:\,u,v\in\mathbb{N}\right\}\subseteq\mathbb{L}(\overline{\mathcal{M}}).$ (6)
Before we study the general structure of $\mathbb{L}(\overline{\mathcal{M}})$ and show that it can generated by a single element $(s_{0},t_{0})$ in the above sense, we consider all non-isomorphic examples in $\operatorname{PG}(3-1,2)$ (ignoring the constraint $s\geq t$).
Example 5.3.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(2,2)$ uniquely characterized by $\mathcal{M}(L)=s\in\mathbb{N}$ for some line $L$ and $\mathcal{M}(L^{\prime})=t\in\mathbb{N}$ for all other lines $L^{\prime}\neq L$. For each point $P\in L$ we have
$\mathcal{M}(P)=\frac{s+2t}{3}-\frac{4t}{6}=\frac{s}{3}$ (7)
and for each point $Q\notin L$ we have
$\mathcal{M}(Q)=\frac{3t}{3}-\frac{s+3t}{6}=\frac{3t-s}{6}.$ (8)
Since $\mathcal{M}(P),\mathcal{M}(Q)\in\mathbb{N}$ we set $x:=\mathcal{M}(P)=\tfrac{s}{3}$ and $y:=\mathcal{M}(Q)=\tfrac{3t-s}{6}$, so that $s=3x$ and $t=2y+x$. With this we have $n=3x+4y$, $\gamma=\max\{x,y\}$, and $s-t=2(x-y)$. If $x\geq y$, then we can write $\mathcal{M}=y\cdot\chi_{E}+(x-y)\cdot\chi_{L}$. If $x\leq y$, then we can write $\mathcal{M}=y\cdot\chi_{E}-(y-x)\cdot\chi_{L}$.
For Example 5.3 the set of all feasible $(s,t)$-pairs assuming $s\geq t$ is given by $\left\langle(3,1)\right\rangle_{\mathbb{N}}+\left\langle(3,3)\right\rangle_{\mathbb{N}}$. If we assume $t\geq s$, then the set of feasible $(s,t)$-pairs is given by $\left\langle(0,2)\right\rangle_{\mathbb{N}}+\left\langle(3,3)\right\rangle_{\mathbb{N}}$. The vector $(0,2)$ can be computed from $(3,1)$ by computing a suitable complement.
Due to Lemma 4.3 we can always assume the existence of a point of multiplicity 00 as a normalization. So, in Example 5.3 we may assume $x=0$ or $y=0$, so that $\mathcal{M}=y\cdot\chi_{E}-y\cdot\chi_{L}$ or $\mathcal{M}=x\cdot\chi_{L}$.
Due to Lemma 4.9 we can always assume that the greatest common divisor of all point multiplicities is $1$ as a normalization (excluding the degenerated case of an empty multiset of points). Applying both normalizations to the multisets of points in Example 5.3 leaves the two possibilities $\chi_{L}$ and $\chi_{E}-\chi_{L}$, i.e., point sets.
Due to Lemma 4.4 we always can assume #$\#\mathcal{M}\leq\gamma(\mathcal{M})\cdot[k]_{q}/2$. Applying also the third normalization to the multisets of points in Example 5.3 leaves only the possibility $\chi_{L}$, i.e., a subspace construction, see Proposition 4.1, where $s=3$, $t=1$, $n=3$, and $s-t=2$.
Example 5.4.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(2,2)$ uniquely characterized by $\mathcal{M}(L_{1})=\mathcal{M}(L_{2})=s\in\mathbb{N}$ for two different lines $L_{1},L_{2}$ and $\mathcal{M}(L^{\prime})=t\in\mathbb{N}$ for all other lines $L^{\prime}\notin\left\{L_{1},L_{2}\right\}$. For $P:=L_{1}\cap L_{2}$ we have
$\mathcal{M}(P)=\frac{2s+t}{3}-\frac{4t}{6}=\frac{2s-t}{3},$ (9)
for each point $Q\in\left(L_{1}\cup L_{2}\right)\backslash\{P\}$ we have
$\mathcal{M}(Q)=\frac{s+2t}{3}-\frac{s+3t}{6}=\frac{s+t}{6},$ (10)
and for each point $R\notin L_{1}\cup L_{2}$ we have
$\mathcal{M}(R)=\frac{3t}{3}-\frac{2s+2t}{6}=\frac{2t-s}{3}.$ (11)
Since $\mathcal{M}(Q),\mathcal{M}(R)\in\mathbb{N}$ we set $x:=\mathcal{M}(Q)=\tfrac{s+t}{6}$ and $y:=\mathcal{M}(R)=\tfrac{2t-s}{3}$, so that $s=4x-y$ and $t=2x+y$. With this we have $n=6x+7y$ and $s-t=2(x-y)$. Of course we need to have $y\leq 2x$ so that $\mathcal{M}(P)\geq 0$, which implies $s\geq 0$.
  • $\mathcal{M}(P)=0$: $y=2x$, so that $\mathcal{M}(P)=0$, $\mathcal{M}(Q)=x$, $\mathcal{M}(R)=2x$, and the greatest common divisor of $\mathcal{M}(P)$, $\mathcal{M}(Q)$, and $\mathcal{M}(R)$ is equal to $x$. Thus, we can assume $x=1$, $y=2$, so that $s=2$, $t=4$, $n=8$, $\gamma=2$, $t-s=2$, and $\mathcal{M}=2\chi_{E}-\chi_{L_{1}}-\chi_{L_{2}}$ for two different lines $L_{1},L_{2}$.
  • $\mathcal{M}(Q)=0$: $x=0$, so that also $y=0$ and $\mathcal{M}$ is the empty multiset of points.
  • $\mathcal{M}(R)=0$: $y=0$, $\mathcal{M}(P)=2x$, $\mathcal{M}(Q)=x$, so that $\gcd(\mathcal{M}(P),\mathcal{M}(Q),\mathcal{M}(R))=x$ and we can assume $x=1$. With this we have $s=4$, $t=2$, $n=6$, $\gamma=2$, $s-t=2$, and $\mathcal{M}=\chi_{L_{1}}+\chi_{L_{2}}$ for two different lines $L_{1},L_{2}$.
So, Example 5.4 can be explained by the construction in Proposition 4.5.
Example 5.5.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(2,2)$ uniquely characterized by $\mathcal{M}(L_{1})=\mathcal{M}(L_{2})=\mathcal{M}(L_{3})=s\in\mathbb{N}$ for three different lines $L_{1},L_{2},L_{3}$ with a common intersection point $P=L_{1}\cap L_{2}\cap L_{3}$ and $\mathcal{M}(L^{\prime})=t\in\mathbb{N}$ for all other lines. We have
$\mathcal{M}(P)=\frac{3s}{3}-\frac{4t}{6}=s-\frac{2t}{3}$ (12)
and
$\mathcal{M}(Q)=\frac{s+2t}{3}-\frac{2s+2t}{6}=\frac{t}{3}$ (13)
for all points $Q\neq P$. Since $\mathcal{M}(P),\mathcal{M}(Q)\in\mathbb{N}$ we set $x:=\mathcal{M}(P)=s-\tfrac{2t}{3}$ and $y:=\mathcal{M}(Q)=\tfrac{t}{3}$, so that $s=x+2y$ and $t=3y$. With this we have $n=x+6y$ and $s-t=x-y$.
  • $\mathcal{M}(P)=0$: $x=0$, so that we can assume $y=1$, which implies $s=2$, $t=3$, $\gamma=1$, $n=6$, $t-s=1$, and $\mathcal{M}=\chi_{E}-\chi_{P}$ for some point $P$.
  • $\mathcal{M}(Q)=0$: $y=0$, so that we can assume $x=1$, which implies $s=1$, $t=0$, $\gamma=1$, $n=1$, $s-t=1$, and $\mathcal{M}=\chi_{P}$ for some point $P$.
So, also Example 5.5 can be explained by the subspace construction in Proposition 4.1.
Example 5.6.
Let $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(2,2)$ uniquely characterized by $\mathcal{M}(L_{1})=\mathcal{M}(L_{2})=\mathcal{M}(L_{3})=s\in\mathbb{N}$ for three different lines $L_{1},L_{2},L_{3}$ without a common intersection point,i.e. $L_{1}\cap L_{2}\cap L_{3}=\emptyset$, and $\mathcal{M}(L^{\prime})=t\in\mathbb{N}$ for all other lines. For each point $P$ that is contained on exactly two lines $L_{i}$ we have
$\mathcal{M}(P)=\frac{2s+t}{3}-\frac{s+3t}{6}=\frac{3s-t}{6},$ (14)
for each point $Q$ that is contained on exactly one line $L_{i}$ we have
$\mathcal{M}(Q)=\frac{s+2t}{3}-\frac{2s+2t}{6}=\frac{t}{3},$ (15)
and for the unique point $R$ that is contained on none of the lines $L_{i}$ we have
$\mathcal{M}(R)=\frac{3t}{3}-\frac{3s+t}{6}=\frac{5t-3s}{6}.$ (16)
Since $\mathcal{M}(P),\mathcal{M}(Q)\in\mathbb{N}$ we set $x:=\mathcal{M}(P)=\tfrac{3s-t}{6}$ and $y:=\mathcal{M}(Q)=\tfrac{t}{3}$, so that $s=2x+y$ and $t=3y$. With this we have $n=2x+5y$ and $s-t=2(x-y)$.
  • $\mathcal{M}(P)=0$: $x=0$, so that we can assume $y=1$, which implies $s=1$, $t=3$, $t-s=2$, $\gamma=2$, $n=5$, and $\mathcal{M}=\chi_{L}+2\chi_{P}$ for some line $L$ and some point $P\notin L$.
  • $\mathcal{M}(Q)=0$: $y=0$, so that $x=0$ and $\mathcal{M}$ is the empty multiset of points.
  • $\mathcal{M}(R)=0$: $x=2y$, so that we can assume $y=1$, which implies $x=2$, $s=4$, $t=6$, $t-s=2$, $\gamma=2$, $n=9$, and the $2$-complement of $\mathcal{M}$ equals $\mathcal{M}=\chi_{L}+2\chi_{P}$ for some line $L$ and some point $P\notin L$, see the case $\mathcal{M}(P)=0$.
So, also Example 5.6 can be explained by the construction in Proposition 4.5.
In Examples 5.35.6 we have considered all cases of $1\leq\#\overline{\mathcal{M}}\leq 3$ up to symmetry. The cases #¯$\#\overline{\mathcal{M}}\in\{0,7\}$ give one-character multisets. By considering the complement $\mathcal{M}^{\prime}=\chi_{V}-\overline{\mathcal{M}}$ we see that examples for $4\leq\#\overline{\mathcal{M}}\leq 6$ do not give something new. Since the dimension of the ambient space is odd, we cannot apply the construction in Proposition 4.6.
Now let us consider the general case. Given the set $\overline{\mathcal{M}}$ of hyperplanes with multiplicity $s$ we get an explicit expression for the multiplicity $\mathcal{M}(P)$ of every point $P\in\mathcal{P}$ depending on the two unknown hyperplane multiplicities $s$ and $t$.
Lemma 5.7.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$, where $k\geq 3$, and $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}(H)=s$ if ¯$\overline{\mathcal{M}}(H^{\perp})=1$ and $\mathcal{M}(H)=t$ if ¯$\overline{\mathcal{M}}(H^{\perp})=0$ for all hyperplanes $H\in\mathcal{H}$. Denoting the number of hyperplanes $H\ni P$ with $\mathcal{M}(H)=s$ by $\varphi(P)$ and setting $r:=\#\overline{\mathcal{M}}$, $\Delta:=s-t\in\mathbb{Z}$, we have
$\mathcal{M}(P)=\frac{t+\Delta\cdot\varphi(P)}{[k-1]_{q}}-\frac{\Delta}{q^{k-2}}\cdot\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-\varphi(P)\right).$ (17)
  • Proof. Counting gives that $[k-1]_{q}-\varphi(P)$ hyperplanes through $P$ have multiplicity $t$, from the $q^{k-1}$ hyperplanes not containing $P$ exactly $r-\varphi(P)$ have multiplicity $\mathcal{M}(H)=s$ and $q^{k-1}-r+\varphi(P)$ have multiplicity $\mathcal{M}(H)=t$. With this we can use Lemma 3.1 to compute
    $\mathcal{M}(P)$ $=$ $\sum_{H\in\mathcal{H}\,:\,P\in H}\frac{1}{[k-1]_{q}}\cdot\mathcal{M}(H)+\sum_{H\in\mathcal{H}\,:\,P\notin H}\frac{1}{q^{k-1}}\cdot\left(\frac{1}{[k-1]_{q}}-1\right)\cdot\mathcal{M}(H)$
    $=$ $\sum_{H\in\mathcal{H}\,:\,P\in H}\frac{1}{[k-1]_{q}}\cdot\mathcal{M}(H)-\sum_{H\in\mathcal{H}\,:\,P\notin H}\frac{1}{q^{k-1}}\cdot\frac{q[k-2]_{q}}{[k-1]_{q}}\cdot\mathcal{M}(H)$
    $=$ $t+\frac{\Delta}{[k-1]_{q}}\cdot\varphi(P)\,\,-\,\,\,\frac{q[k-2]_{q}}{[k-1]_{q}}\cdot t-\frac{\Delta}{q^{k-1}}\cdot\frac{q[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-\varphi(P)\right)$
    $=$ $\frac{t+\Delta\cdot\varphi(P)}{[k-1]_{q}}-\frac{\Delta}{q^{k-2}}\cdot\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-\varphi(P)\right).$
    $\square$
Note that $\varphi(P)=\overline{\mathcal{M}}(P^{\perp})$ for all $P\in\mathcal{P}$.
Lemma 5.8.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$, where $k\geq 3$, and $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}(H)=s$ if ¯$\overline{\mathcal{M}}(H^{\perp})=1$ and $\mathcal{M}(H)=t$ if ¯$\overline{\mathcal{M}}(H^{\perp})=0$ for all hyperplanes $H\in\mathcal{H}$. Denote the number of hyperplanes $H\ni P$ with $\mathcal{M}(H)=s$ by $\varphi(P)$ and uniquely choose $m\in\mathbb{N}$, $\mathcal{I}\subseteq\mathbb{N}$ with $0\in\mathcal{I}$ such that $\left\{\varphi(P)\,:\,P\in\mathcal{P}\right\}=\left\{m+i\,:\,i\in\mathcal{I}\right\}$. If $s>t$ and there exists a point $Q\in\mathcal{P}$ with $\mathcal{M}(Q)=0$, then we have
$t=\frac{\Delta}{q^{k-2}}\cdot[k-2]_{q}\cdot\left(r-m\right)-\Delta\cdot m$ (18)
and
$\mathcal{M}(P)=\frac{\Delta\cdot i}{q^{k-2}}$ (19)
for all points $P\in\mathcal{P}$ where $i:=\varphi(P)-m$, $r:=\#\overline{\mathcal{M}}$, and $\Delta:=s-t\in\mathbb{N}_{\geq 1}$. If $\mathcal{M}$ is non-repetitive, then $\Delta$ divides $q^{k-2}$.
  • Proof. Using $\Delta>0$ we observe that the expression for $\mathcal{M}(P)$ in Equation (17) is increasing in $\varphi(P)$. So, we need to choose a point $Q\in\mathcal{P}$ which minimizes $\varphi(Q)$ to normalize using $\mathcal{M}(Q)=0$ since otherwise we will obtain points with negative multiplicity. So, choosing a point $Q\in\mathcal{P}$ with $\varphi(Q)=m$ we require
    $0=\mathcal{M}(Q)=\frac{t+\Delta\cdot m}{[k-1]_{q}}-\frac{\Delta}{q^{k-1}}\cdot\frac{q[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-m\right),$
    which yields Equation (18). Using $i:=\varphi(P)-m$ and the expression for $t$ we compute
    $\mathcal{M}(P)$ $=$ $\frac{t+\Delta\cdot(m+i)}{[k-1]_{q}}-\frac{\Delta}{q^{k-2}}\cdot\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-m-i\right)$
    $=$ $\frac{\Delta}{q^{k-2}}\cdot\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-m\right)-\frac{\Delta\cdot m}{[k-1]_{q}}+\frac{\Delta\cdot(m+i)}{[k-1]_{q}}-\frac{\Delta}{q^{k-2}}\cdot\frac{[k-2]_{q}}{[k-1]_{q}}\cdot\left(r-m-i\right)$
    $=$ $\frac{\Delta\cdot i}{[k-1]_{q}}+\frac{\Delta\cdot i}{q^{k-2}}\cdot\frac{[k-2]_{q}}{[k-1]_{q}}=\frac{\Delta\cdot i}{q^{k-2}}$
    for all $P\in\mathcal{P}$. Note that if $f>1$ is a divisor of $\Delta$ that is coprime to $q$, then all point multiplicities of $\mathcal{M}$ are divisible by $f$. If $\Delta=q^{k-2}\cdot f$ for an integer $f>1$, then all point multiplicities of $\mathcal{M}$ are divisible by $f$. Thus, we have that $\Delta$ divides $q^{k-2}$. $\square$
Note that ={¯$\mathcal{I}=\left\{\overline{\mathcal{M}}(H)-\overline{\mathcal{M}}(H^{\prime})\,:\,H\in\mathcal{H}\right\}$, where $H^{\prime}\in\mathcal{H}$ is a minimizer of ¯$\overline{\mathcal{M}}(H)$.
Lemma 5.9.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$, where $k\geq 3$ and $\mathcal{M}$ be a multiset of points in $\operatorname{PG}(k-1,q)$ such that $\mathcal{M}(H)=s$ if ¯$\overline{\mathcal{M}}(H^{\perp})=1$ and $\mathcal{M}(H)=t$ if ¯$\overline{\mathcal{M}}(H^{\perp})=0$ for all hyperplanes $H\in\mathcal{H}$. Using the notation from Lemma 5.8 we set
$g$ $=$ $gcd\!\left(\{i\in\mathcal{I}\}\cup\{q^{k-2}\}\right),$ (20)
$\Delta_{0}$ $=$ $q^{k-2}/g,$ (21)
$t_{0}$ $=$ $\frac{1}{g}\cdot[k-2]_{q}\cdot\left(r-m\right)-\Delta_{0}\cdot m,\text{ and}$ (22)
$s_{0}$ $=$ $t+\Delta_{0}.$ (23)
If $s>t$, then we have
$\mathbb{L}(\overline{\mathcal{M}})=\left\langle(s_{0},t_{0})\right\rangle_{\mathbb{N}}+\left\langle([k-1]_{q},[k-1]_{q})\right\rangle_{\mathbb{N}}.$
  • Proof. Setting $\mu=\mu(\mathcal{M})\in\mathbb{N}$ we have that $\mathcal{M}^{\prime}:=\mathcal{M}-\mu\cdot\chi_{V}$ is a two-character multiset corresponding to $(s^{\prime},t^{\prime}):=(s-\mu[k-1]_{q},t-\mu[k-1]_{q})\in\mathbb{L}(\overline{\mathcal{M}})$ and there exists a point $Q\in\mathcal{P}$ with $\mathcal{M}^{\prime}(Q)=0$. Clearly, we have $(s^{\prime},t^{\prime})\in\mathbb{N}^{2}$ and $s^{\prime}>t^{\prime}$. From Lemma 5.8 we conclude the existence of an integer $\Delta^{\prime}\in\mathbb{N}_{\geq 1}$ such that $t^{\prime}=\frac{\Delta^{\prime}}{q^{k-2}}\cdot[k-2]_{q}\cdot\left(r-m\right)-\Delta^{\prime}\cdot m$, $s^{\prime}=t^{\prime}+\Delta^{\prime}$, and $\mathcal{M}^{\prime}(P)=\frac{\Delta^{\prime}\cdot i}{q^{k-2}}$ for all $P\in\mathcal{P}$. Since $\mathcal{M}^{\prime}(P)\in\mathbb{N}$ for all $P\in\mathcal{P}$ we have that $q^{k-2}$ divides $\Delta^{\prime}\cdot g$, so that $\Delta_{0}\in\mathbb{N}$ divides $\Delta^{\prime}$. For $f:=\Delta^{\prime}/\Delta_{0}\in\mathbb{N}_{\geq 1}$ we observe that $\mathcal{M}^{\prime}(P)$ is divisible by $f$ and we set ′′$\mathcal{M}^{\prime\prime}:=\tfrac{1}{f}\cdot\mathcal{M}^{\prime}$. With this, we can check that $\mathcal{M}^{\prime\prime}$ is a two-character multiset corresponding to $(s_{0},t_{0})\in\mathbb{L}(\overline{\mathcal{M}})$. $\square$
Note that it is not necessary to explicitly check $t_{0}\in\mathbb{N}$ since ′′$\mathcal{M}^{\prime\prime}(P)\in\mathbb{N}$ is sufficient to this end.
Before we consider the problem whether $\mathbb{L}(\overline{\mathcal{M}})\subseteq\mathbb{N}^{2}$ contains an element $(s,t)$ with $s>t$ we treat the so far excluded case $k=2$ separately.
Lemma 5.10.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(1,q)$. Then, we have
$\mathbb{L}(\overline{\mathcal{M}})=\left\langle(s_{0},0)\right\rangle_{\mathbb{N}}+\left\langle(q+1,q+1)\right\rangle_{\mathbb{N}},$
where $s_{0}=0$ if #¯$\#\overline{\mathcal{M}}\in\{0,q+1\}$ and $s_{0}=1$ otherwise.
  • Proof. If #¯$\#\overline{\mathcal{M}}\in\{0,q+1\}$, then a two-character multiset $\mathcal{M}$ corresponding to $(s,t)\in\overline{\mathcal{M}}$ actually is a one-character multiset and there exist some integer $x\in\mathbb{N}$ such that $\mathcal{M}=x\cdot\chi_{v}$.
    Otherwise we observe that in $\operatorname{PG}(1,q)$ points and hyperplanes coincide and the image of $\overline{\mathcal{M}}$ is $\{0,1\}$. Note that we have $\mathcal{M}=t\cdot\chi_{V}+\sum_{P\in\mathcal{P}}(s-t)\cdot\overline{\mathcal{M}}(P)\cdot\chi_{P}$ for each two-character multiset $\mathcal{M}$ corresponding to $(s,t)\in\mathbb{L}(\overline{\mathcal{M}})$ by definition. We can easily check $(s,t)\in\left\langle(1,0)\right\rangle_{\mathbb{N}}+\left\langle(q+1,q+1)\right\rangle_{\mathbb{N}}$. The proof is completed by choosing $s=1$ and $t=0$ in our representation of $\mathcal{M}$. $\square$
Theorem 5.11.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$ with #¯$\#\overline{\mathcal{M}}\notin\left\{0,[k]_{q}\right\}$, where $k\geq 2$. Then
$\mathcal{M}:=\sum_{H\in\mathcal{H}}\overline{\mathcal{M}}(H^{\perp})\cdot\chi_{H}$ (24)
is a two-character multiset corresponding to $(s,t)\in\mathbb{L}(\overline{\mathcal{M}})$ with $n=\#\mathcal{M}=r[k-1]_{q}$, where $r:=\#\overline{\mathcal{M}}$, $t=r[k-2]_{q}$, and $s=r[k-2]_{q}+q^{k-2}$. Setting $\mu:=\mu(\mathcal{M})$ and $g:=\operatorname{gcd}(\left\{\mathcal{M}(P)-\mu\,:\,P\in\mathcal{P}\right\})$ the function
$\mathcal{M}^{\prime}:=\frac{1}{g}\cdot\left(-\mu\cdot\chi_{V}+\sum_{H\in\mathcal{H}}\overline{\mathcal{M}}(H^{\perp})\cdot\chi_{H}\right)=\frac{1}{g}\cdot\left(\mathcal{M}-\mu\cdot\chi_{V}\right)$ (25)
is a two-character multiset corresponding to $(s_{0},t_{0})\in\mathbb{L}(\overline{\mathcal{M}})$ with $n^{\prime}=\#\mathcal{M}^{\prime}=\tfrac{1}{q}\cdot\left(r[k-1]_{q}-\mu[k]_{q}\right)$, where $r:=\#\overline{\mathcal{M}}$, $t_{0}=\tfrac{1}{g}\cdot\left(r[k-2]_{q}-\mu[k-1]_{q}\right)$, and $s_{0}=\tfrac{1}{g}\cdot\left(r[k-2]_{q}-\mu[k-1]_{q}+q^{k-2}\right)$, and $g$ divides $q^{k-2}$. Moreover, we have
$\mathbb{L}(\overline{\mathcal{M}})=\left\langle(s_{0},t_{0})\right\rangle_{\mathbb{N}}+\left\langle([k-1]_{q},[k-1]_{q})\right\rangle_{\mathbb{N}},$ (26)
  • Proof. We can easily check $\mathcal{M}(H)=r[k-2]_{q}=t$ if $\mathcal{M}(H^{\perp})=0$ and $\mathcal{M}(H)=r[k-2]_{q}+q^{k-2}=s$ if $\mathcal{M}(H^{\perp})=1$ for all $H\in\mathcal{H}$ as well as #$\#\mathcal{M}=r[k-1]_{q}$ directly from the definition of $\mathcal{M}$. Using Lemma 4.3 and Lemma 4.9 we conclude that $\mathcal{M}^{\prime}$ is a two-character multiset with the stated parameters.
    For $k=2$ Lemma 5.10 our last statement. For $k\geq 3$ we can apply Lemma 5.8 to conclude $g=\operatorname{gcd}(\{i\in\mathcal{I}\})$ and use the proof of Lemma 5.9 to conclude our last statement. Since $s,t\in\mathbb{N}$ and $s>t$ we have that $g$ divides $g(s-t)=q^{k-2}$. $\square$
Using the notation from Lemma 5.8 applied to to the multiset of points $\mathcal{M}-\mu\cdot\chi_{V}$ from Theorem 5.11 we observe #$\#\mathcal{I}\geq 2$ for #¯$\#\overline{\mathcal{M}}\notin\left\{0,[k]_{q}\right\}$. Using the fact that $g:=\operatorname{gcd}(\left\{\mathcal{M}(P)-\mu\,:\,P\in\mathcal{P}\right\})$, that $g$ divides $q^{k-2}$, and Equation (19) we conclude
$g=gcd\!\left(\{i\in\mathcal{I}\}\right)=gcd\!\left(\left\{\overline{\mathcal{M}}(H)-\overline{\mathcal{M}}(H^{\prime})\,:\,H\in\mathcal{H}\right\}\right),$ (27)
where $H^{\prime}\in\mathcal{H}$ is a minimizer of ¯$\overline{\mathcal{M}}(H)$.
Using the classification of one-character multisets we conclude from Theorem 5.11:
Corollary 5.12.
Let $\overline{\mathcal{M}}$ be a set of points in $\operatorname{PG}(k-1,q)$, where $k\geq 2$. Then, there exist $\left(s_{0},t_{0}\right)\in\mathbb{N}^{2}$ such that $\mathbb{L}(\overline{\mathcal{M}})=\left\langle(s_{0},t_{0})\right\rangle_{\mathbb{N}}+\left\langle([k-1]_{q},[k-1]_{q})\right\rangle_{\mathbb{N}}$.
Theorem 5.13.
Let $-~\widetilde{\mathcal{M}}$ be a two-character multiset in $\operatorname{PG}(k-1,q)$, where $k\geq 2$. Then, there exist unique $u,v\in\mathbb{N}$ such that ~$\widetilde{\mathcal{M}}=u\cdot\mathcal{M}^{\prime}+v\cdot\chi_{V}$, where $\mathcal{M}^{\prime}$ is given by Equation (25).
  • Proof. Let $s>t$ be the two hyperplane multiplicities of $\widetilde{\mathcal{M}}$. With this define $\overline{\mathcal{M}}$ such that ¯$\overline{\mathcal{M}}(H^{\perp})=1$ if ~$\widetilde{\mathcal{M}}(H)=s$ and ¯$\overline{\mathcal{M}}(H^{\perp})=0$ if ~$\widetilde{\mathcal{M}}(H)=t$ for all $H\in\mathcal{H}$. So, $(s,t)\in\mathbb{L}(\overline{\mathcal{M}})$ and Theorem 5.11 yields the existence of $u,v\in\mathbb{N}$ with $(s,t)=u\cdot(s_{0},t_{0})+v\cdot([k-1]_{q},[k-1]_{q})$, where $s_{0}$, $t_{0}$ are as in Theorem 5.11. From Lemma 3.1 we then conclude ~$\widetilde{\mathcal{M}}=u\cdot\mathcal{M}^{\prime}+v\cdot\chi_{V}$. Note that $\mu(\mathcal{M}^{\prime})$ and $\mu(\chi_{V})=1$ imply $\mu(\widetilde{M})=v$, so that $u$ can be computed from $\gamma(\widetilde{\mathcal{M}})=u\cdot\gamma(\mathcal{M}^{\prime})+v$. $\square$
Note that for a one-character multiset $\widetilde{\mathcal{M}}$ there exists a unique $v\in\mathbb{N}$ such that ~$\widetilde{\mathcal{M}}=v\cdot\chi_{V}$. Given a set of points $\overline{\mathcal{M}}$ we call $\mathcal{M}^{\prime}$ the canonical representant of the set of two-character multisets $\mathcal{M}$ corresponding to $(s,t)\in\mathbb{L}(\overline{\mathcal{M}})$. If $\mathcal{M}=\mathcal{M}^{\prime}$ we just say that $\mathcal{M}$ is the canonical two-character multiset.
Theorem 5.14.
Let $w_{1}<w_{2}$ be the non-zero weights of a non-repetitive $[n,k]_{q}$ two-weight code $C$ without full support. Then, there exist integers $f$ and $u$ such that $w_{1}=up^{f}$ and $w_{2}=(u+1)p^{f}$, where $p$ is the characteristic of the underlying field $\mathbb{F}_{q}$.
  • Proof. Let $\mathcal{M}$ be the two-character multiset in $\operatorname{PG}(k-1,q)$ corresponding to $C$. Choose unique $u,v\in\mathbb{N}$ such that $\mathcal{M}=u\cdot\mathcal{M}^{\prime}+v\cdot\chi_{V}$ as in Theorem 5.13. Since $C$ does not have full support, we have $v=0$ and since $C$ is non-repetitive we have $u=1$. With this we can use Theorem 5.11 to compute
    $w_{1}=n-s=\tfrac{1}{g}\cdot\left(r\cdot q^{k-2}-\mu\cdot q^{k-1}-q^{k-2}\right)=(r-q\mu-1)\cdot p^{f}$ (28)
    and
    $w_{2}=n-t=\tfrac{1}{g}\cdot\left(r\cdot q^{k-2}-\mu\cdot q^{k-1}\right)=(r-q\mu)\cdot p^{f},$ (29)
    where $f$ is chosen such that $\tfrac{q^{k-2}}{g}=p^{f}$. I.e., we can choose $u=r-q\mu-1$. $\square$
We have seen in Equation (27) that we can compute the parameter $g$ directly from the set of points $\overline{\mathcal{M}}$. If we additionally assume that $\overline{\mathcal{M}}$ is spanning, then we can consider the corresponding projective $[n,k]_{q}$-code $\overline{C}$, where $n=\#\overline{\mathcal{M}}$. [^3] If $\overline{\mathcal{M}}$ is not spanning then we can consider the lowerdimensional subspace spanned by $\operatorname{supp}(\overline{\mathcal{M}})$. Note that we have ¯$\overline{\mathcal{M}}(H)\equiv m\pmod{g}$ for all $H\in\mathcal{H}$ and that $g$ is maximal with this property. If $m\equiv n\pmod{g}$, then $g$ would simply be the maximal divisibility constant of the weights of $\overline{C}$. From [29, Theorem 7] or [40, Theorem 3] we can conclude $m\equiv n\pmod{g}$. Thus, we have
$g=\operatorname{gcd}\!\left(\left\{\operatorname{wt}(c)\,:\,c\in\overline{C}\right\}\right).$ (30)
The argument may also be based on the following lemma (using the fact that $\overline{C}$ is projective):
Lemma 5.15.
Let $C$ be an $[n,k]_{q}$-code of full length such that we have $\operatorname{wt}(c)\equiv m\pmod{\Delta}$ for all non-zero codewords $c\in C$. If $\Delta$ is a power of the characteristic of the underlying field $\mathbb{F}_{q}$, then we have $m\equiv 0\pmod{\min\{\Delta,q\}}$. Moreover, if additionally $q$ divides $\Delta$ and $k\geq 2$, then the non-zero weights in each residual code are congruent to $m/q$ modulo $\Delta/q$.
  • Proof. Let $\mathcal{M}$ be the multiset of points in $\operatorname{PG}(k-1,q)$ corresponding to $C$. For each hyperplane $H$ we have $n-\mathcal{M}(H)\equiv m\pmod{\Delta}$, which is equivalent to $\mathcal{M}(H)\equiv n-m\pmod{\Delta}$. The weight of a non-zero codeword in a residual code is given by a subspace $K$ of codimension $2$ and a hyperplane $H$ with $K\leq H$. With this, the weight is given by $\mathcal{M}(H)-\mathcal{M}(K)\equiv n-m-\mathcal{M}(K)\pmod{\Delta}$. Counting the hyperplane multiplicities of the $q+1$ hyperplanes that contain $K$ yields
    $\sum_{H\in\mathcal{H}\,:\,K\leq H}\mathcal{M}(H)=\#\mathcal{M}+q\cdot\mathcal{M}(K)=\#\mathcal{M}+q\cdot\mathcal{M}(K)$ (31)
    and
    $\sum_{H\in\mathcal{H}\,:\,K\leq H}\mathcal{M}(H)\equiv(q+1)(n-m)\pmod{\Delta},$ (32)
    so that
    $m\equiv q\cdot\left(n-m-\mathcal{M}(K)\right)\pmod{\Delta}.$ (33)
    $\square$
Given Equation (30) we might be interested in projective divisible codes (with a large divisibility constant). For enumerations for the binary case we refer to [27] and for a more general survey we refer to e.g. [34]. Note that the only point sets $\mathcal{M}$ in $\operatorname{PG}(k-1,q)$ that are $q^{k-1}$-divisible are given by #$\#\mathcal{M}\in\left\{0,[k]_{q}\right\}$, i.e., the empty and the full set. All other point sets are at most $q^{k-2}$-divisible, as implied by Theorem 5.11.
6. Enumeration of two-character multisets in $\operatorname{PG}(k-1,q)$ for small parameters
Since all two-character multisets in $\operatorname{PG}(1,q)$ can be parameterized as $\mathcal{M}=b\cdot\chi_{V}+\sum_{P\in\mathcal{P}}(a-b)\cdot\overline{\mathcal{M}}(P)\cdot\chi_{P}$ for integers $a>b\geq 0$ and a set of points $\overline{\mathcal{M}}$ in $\operatorname{PG}(k-1,q)$, see Lemma 5.10 and its proof, we assume $k\geq 3$ in the following. Due to Theorem 5.13 every two-character multiset in $\operatorname{PG}(k-1,q)$ can be written as $u\cdot\mathcal{M}^{\prime}+v\cdot\chi_{V}$, where $u,v\in\mathbb{N}$ and $\mathcal{M}^{\prime}$ is characterized in Theorem 5.11. So, we further restrict out considerations on canonical two-character multisets where we have $u=1$ and $v=0$. For $k=2$ all canonical two-character multisets in $\operatorname{PG}(k-1,q)$ are indeed sets of points and given by the construction in Proposition 4.6 (with $r=1$).
It can be easily checked that two isomorphic sets of points in $\operatorname{PG}(k-1,q)$ yield isomorphic canonical two-character multisets $\mathcal{M}^{\prime}$. So, for the full enumeration of canonical two-character multisets in $\operatorname{PG}(k-1,q)$ we just need to loop over all non-isomorphic sets of points $\overline{\mathcal{M}}$ in $\operatorname{PG}(k-1,q)$ and use Theorem 5.11 to determine $\mathcal{M}$, $\mathcal{M}^{\prime}$, and their parameters. We remark that the numbers of non-isomorphic projective codes per length, dimension, and field size are e.g. listed in [1, Tables 6.10–6.12] (for small parameters). For the binary case and dimensions at most six some additional data can be found in [4]. Here we utilize the software package LinCode [6] to enumerate these codes.
$g$ $\mu$ $r$ $n$ $\gamma$ $s$ $t$ $s_{0}$ $t_{0}$ $n^{\prime}$ $\gamma^{\prime}$
2 1 3 9 3 5 3 1 0 1 1
1 0 1 3 1 3 1 3 1 3 1
1 2 6 18 3 8 6 2 0 4 1
2 0 4 12 2 6 4 3 2 6 1
1 1 4 12 3 6 4 3 1 5 2
1 0 2 6 2 4 2 4 2 6 2
1 1 5 15 3 7 5 4 2 8 2
1 0 3 9 2 5 3 5 3 9 2
Table 1. Feasible parameters for canonical two-character multisets in $\operatorname{PG}(2,2)$.
In Table 1 and in Table 2 we list the feasible parameters for canonical two-character multisets in $\operatorname{PG}(2,2)$ and in $\operatorname{PG}(3,2)$, respectively, where $n^{\prime}:=\#\mathcal{M}^{\prime}$ and $\gamma^{\prime}:=\gamma(\mathcal{M}^{\prime})$. For $\operatorname{PG}(2,2)$ we can also state more direct constructions:
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(1,1,0,1)$: characteristic function of a point (not spanning)
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(3,3,1,1)$: characteristic function of a line (not spanning)
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(4,2,0,1)$: complement of the characteristic function of a line
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(6,3,2,1)$: complement of the characteristic function of a point
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(5,3,1,2)$: $\chi_{L}+2\chi_{P}$ for a line $L$ and a point $P$ with $P\notin L$
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(6,4,2,2)$: $\chi_{L}+\chi_{L}^{\prime}$ for two different lines $L$ and $L^{\prime}$
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(8,4,2,2)$: $\chi_{V}-\chi_{L}-\chi_{L}^{\prime}$ for two different lines $L$ and $L^{\prime}$
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(9,5,3,2)$: $2\chi_{V}-\chi_{L}+-\chi_{P}$ for a line $L$ and a point $P$ with $P\notin L$
$g$ $\mu$ $r$ $n$ $\gamma$ $s$ $t$ $s_{0}$ $t_{0}$ $n^{\prime}$ $\gamma^{\prime}$
4 3 7 49 7 25 21 1 0 1 1
2 1 3 21 3 13 9 3 1 3 1
2 4 10 70 6 34 30 3 1 5 1
2 2 6 42 4 22 18 4 2 6 1
1 0 1 7 1 7 3 7 3 7 1
1 6 14 98 7 46 42 4 0 8 1
2 3 9 63 5 31 27 5 3 9 1
2 1 5 35 3 19 15 6 4 10 1
2 4 12 84 6 40 36 6 4 12 1
4 0 8 56 4 28 24 7 6 14 1
2 2 8 56 6 28 24 7 5 13 2
1 0 2 14 2 10 6 10 6 14 2
2 0 4 28 4 16 12 8 6 14 2
1 5 13 91 7 43 39 8 4 16 2
2 3 11 77 7 37 33 8 6 16 2
2 1 7 49 5 25 21 9 7 17 2
1 1 4 28 4 16 12 9 5 13 3
1 4 11 77 7 37 33 9 5 17 3
1 3 9 63 6 31 27 10 6 18 3
1 2 7 49 5 25 21 11 7 19 3
1 1 5 35 4 19 15 12 8 20 3
1 0 3 21 3 13 9 13 9 21 3
1 4 12 84 7 40 36 12 8 24 3
1 3 10 70 6 34 30 13 9 25 3
1 2 8 56 5 28 24 14 10 26 3
1 1 6 42 4 22 18 15 11 27 3
1 0 4 28 3 16 12 16 12 28 3
1 3 11 77 6 37 33 16 12 32 3
1 3 8 56 7 28 24 7 3 11 4
1 2 6 42 6 22 18 8 4 12 4
1 3 9 63 7 31 27 10 6 18 4
1 2 7 49 6 25 21 11 7 19 4
1 1 5 35 5 19 15 12 8 20 4
1 3 10 70 7 34 30 13 9 25 4
1 2 8 56 6 28 24 14 10 26 4
1 1 6 42 5 22 18 15 11 27 4
1 2 9 63 6 31 27 17 13 33 4
1 1 7 49 5 25 21 18 14 34 4
1 0 5 35 4 19 15 19 15 35 4
1 2 10 70 6 34 30 20 16 40 4
1 1 8 56 5 28 24 21 17 41 4
1 0 6 42 4 22 18 22 18 42 4
1 1 9 63 5 31 27 24 20 48 4
1 0 7 49 4 25 21 25 21 49 4
Table 2. Feasible parameters for canonical two-character multisets in $\operatorname{PG}(3,2)$.
Of course, also for $\operatorname{PG}(3,2)$ some of the examples have nicer descriptions:
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(1,1,0,1)$: characteristic function of a point (not spanning)
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(3,3,1,1)$: characteristic function of a line (not spanning)
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(7,7,3,1)$: characteristic function of a plane (not spanning)
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(5,3,1,1)$: projective base; spanning projective $2$-weight code
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(6,4,2,1)$: characteristic function of two disjoint lines; spanning projective $2$-weight code
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(14,10,6,2)$: characteristic function of two different planes
  • $(n^{\prime},s_{0},t_{0},\gamma^{\prime})=(21,13,9,3)$: characteristic function of three planes intersecting in a common point but not a common line
Note that we may restrict our considerations to $r<[k]_{q}/2$ since if $\mathcal{M}^{\prime}$ is the a canonical two-character multiset for a set of points $\overline{\mathcal{M}}$ with #¯$\#\overline{\mathcal{M}}=r$, then the complement of $\mathcal{M}^{\prime}$ is the the a canonical two-character multiset for a set of points which is the complement of $\overline{\mathcal{M}}$ and has cardinality $[k]_{q}-r$.
From the data in Table 1 and Table 2 we can guess the the maximum possible point multiplicity $\gamma(\mathcal{M}^{\prime})$ of $\mathcal{M}^{\prime}$:
Proposition 6.1.
Let $\mathcal{M}$ be a canonical two-character multiset in $\operatorname{PG}(k-1,q)$, where $k\geq 2$. Then, we have $\gamma(\mathcal{M})\leq q^{k-2}$.
  • Proof. Choose a suitable set $\mathcal{H}^{\prime}\subseteq\mathcal{H}$ and $g,\nu\in\mathbb{N}$ such that
    $\mathcal{M}=\frac{1}{g}\cdot\left(\sum_{H\in\mathcal{H}^{\prime}}\chi_{H}\,-\,\mu\cdot\chi_{V}\right).$
    Let $P\in\mathcal{P}$ be a point with $\mathcal{M}(P)=\gamma$ and $Q\in\mathcal{P}$ be a point with $\mathcal{M}(Q)=0$. With this we have $\lambda\geq\left|\left\{H\in\mathcal{H}^{\prime}\,:\,Q\leq H\right\}\right|$. Since $P$ is contained $[k-1]_{q}$ hyperplanes in $\mathcal{H}$ and $\langle P,Q\rangle$ is contained in $[k-2]_{q}$ hyperplanes in $\mathcal{H}$ we have $\mathcal{M}(P)\leq q^{k-2}$. $\square$
We can easily construct an example showing that the stated upper bound is tight. To this end let $P$, $Q$ be two different points in $\operatorname{PG}(k-1,q)$, where $k\geq 3$,and $H^{\prime}$ be an arbitrary hyperplane neither containing $P$ nor $Q$. With this, we choose $\mathcal{H}^{\prime}$ as the set of all $q^{k-2}$ hyperplanes that contain $P$ but do not contain $Q$ and additionally the hyperplane $H^{\prime}$. For the corresponding multiset of points $\mathcal{M}$ we then have $\mathcal{M}(P)=q^{k-2}$ and $\mathcal{M}(Q)=0$, so that $\mu(\mathcal{M})=0$. For an arbitrary point $R\in H^{\prime}$ we have $\mathcal{M}(R)=q^{k-2}-q^{k-3}+1=(q-1)q^{k-3}+1$, so that $\operatorname{gcd}(\mathcal{M}(R),\mathcal{M}(P))=1$ if $k\geq 4$ or $k=3$ and $q\neq 2$. For $(k,q)=(3,2)$ we have already seen examples of canonical two-character multisets with maximum point multiplicity $2$.
In Table 3 and Table 4 we list the feasible parameters for canonical two-character multisets in $\operatorname{PG}(4,2)$ with point multiplicity at most $4$.
$g$ $\mu$ $r$ $n$ $\gamma$ $s$ $t$ $s_{0}$ $t_{0}$ $n^{\prime}$ $\gamma^{\prime}$
8 7 15 225 15 113 105 1 0 1 1
4 3 7 105 7 57 49 3 1 3 1
2 1 3 45 3 29 21 7 3 7 1
1 0 1 15 1 15 7 15 7 15 1
1 14 30 450 15 218 210 8 0 16 1
2 12 28 420 14 204 196 12 8 24 1
4 8 24 360 12 176 168 14 12 28 1
8 0 16 240 8 120 112 15 14 30 1
2 2 6 90 6 50 42 10 6 14 2
2 5 13 195 9 99 91 12 8 20 2
2 3 9 135 7 71 63 13 9 21 2
2 1 5 75 5 43 35 14 10 22 2
2 8 20 300 12 148 140 14 10 26 2
2 6 16 240 10 120 112 15 11 27 2
2 4 12 180 8 92 84 16 12 28 2
2 2 8 120 6 64 56 17 13 29 2
4 0 8 120 8 64 56 16 14 30 2
2 0 4 60 4 36 28 18 14 30 2
1 0 2 30 2 22 14 22 14 30 2
4 7 23 345 15 169 161 16 14 32 2
2 11 27 405 15 197 189 16 12 32 2
1 13 29 435 15 211 203 16 8 32 2
2 9 23 345 13 169 161 17 13 33 2
2 7 19 285 11 141 133 18 14 34 2
2 5 15 225 9 113 105 19 15 35 2
2 3 11 165 7 85 77 20 16 36 2
2 10 26 390 14 190 182 20 16 40 2
2 8 22 330 12 162 154 21 17 41 2
2 6 18 270 10 134 126 22 18 42 2
2 9 25 375 13 183 175 24 20 48 2
2 4 10 150 10 78 70 9 5 13 3
2 9 21 315 15 155 147 10 6 18 3
2 3 9 135 9 71 63 13 9 21 3
2 6 16 240 12 120 112 15 11 27 3
2 4 12 180 10 92 84 16 12 28 3
2 2 8 120 8 64 56 17 13 29 3
1 1 4 60 4 36 28 21 13 29 3
2 7 19 285 13 141 133 18 14 34 3
2 5 15 225 11 113 105 19 15 35 3
2 3 11 165 9 85 77 20 16 36 3
2 1 7 105 7 57 49 21 17 37 3
2 6 18 270 12 134 126 22 18 42 3
2 4 14 210 10 106 98 23 19 43 3
2 2 10 150 8 78 70 24 20 44 3
1 0 3 45 3 29 21 29 21 45 3
1 12 28 420 15 204 196 24 16 48 3
2 7 21 315 13 155 147 25 21 49 3
Table 3. Feasible parameters for canonical two-character multisets in $\operatorname{PG}(4,2)$ with $\gamma^{\prime}\leq 4$ – part 1.
$g$ $\mu$ $r$ $n$ $\gamma$ $s$ $t$ $s_{0}$ $t_{0}$ $n^{\prime}$ $\gamma^{\prime}$
2 5 17 255 11 127 119 26 22 50 3
2 3 13 195 9 99 91 27 23 51 3
2 8 24 360 14 176 168 28 24 56 3
2 6 20 300 12 148 140 29 25 57 3
2 4 16 240 10 120 112 30 26 58 3
2 2 12 180 8 92 84 31 27 59 3
1 11 27 405 14 197 189 32 24 64 3
2 7 23 345 13 169 161 32 28 64 3
2 5 19 285 11 141 133 33 29 65 3
2 3 15 225 9 113 105 34 30 66 3
2 6 22 330 12 162 154 36 32 72 3
2 0 10 150 6 78 70 39 35 75 3
2 5 21 315 11 155 147 40 36 80 3
2 4 12 180 12 92 84 16 12 28 4
1 2 6 90 6 50 42 20 12 28 4
2 3 11 165 11 85 77 20 16 36 4
2 4 14 210 12 106 98 23 19 43 4
1 2 7 105 6 57 49 27 19 43 4
1 1 5 75 5 43 35 28 20 44 4
1 11 26 390 15 190 182 25 17 49 4
1 10 24 360 14 176 168 26 18 50 4
1 9 22 330 13 162 154 27 19 51 4
1 8 20 300 12 148 140 28 20 52 4
1 7 18 270 11 134 126 29 21 53 4
1 5 14 210 9 106 98 31 23 55 4
2 6 20 300 14 148 140 29 25 57 4
1 3 10 150 7 78 70 33 25 57 4
2 4 16 240 12 120 112 30 26 58 4
1 2 8 120 6 64 56 34 26 58 4
1 1 6 90 5 50 42 35 27 59 4
1 0 4 60 4 36 28 36 28 60 4
1 11 27 405 15 197 189 32 24 64 4
1 10 25 375 14 183 175 33 25 65 4
2 3 15 225 11 113 105 34 30 66 4
1 9 23 345 13 169 161 34 26 66 4
2 1 11 165 9 85 77 35 31 67 4
1 8 21 315 12 155 147 35 27 67 4
1 6 17 255 10 127 119 37 29 69 4
1 4 13 195 8 99 91 39 31 71 4
1 3 11 165 7 85 77 40 32 72 4
1 2 9 135 6 71 63 41 33 73 4
1 1 7 105 5 57 49 42 34 74 4
1 0 5 75 4 43 35 43 35 75 4
1 10 26 390 14 190 182 40 32 80 4
2 3 17 255 11 127 119 41 37 81 4
1 9 24 360 13 176 168 41 33 81 4
2 4 20 300 12 148 140 44 40 88 4
2 3 19 285 11 141 133 48 44 96 4
1 9 25 375 13 183 175 48 40 96 4
Table 4. Feasible parameters for canonical two-character multisets in $\operatorname{PG}(4,2)$ with $\gamma^{\prime}\leq 4$ – part 2.
Appendix A Feasible parameters for binary projective two-weight codes
In this appendix we want to utilize our parameterization for the parameters of two-character multisets in order to determine the set of feasible parameters of two-character sets being equivalent to projective two-weight codes.
Starting with the sum of the characteristic functions of $r$ different hyperplanes in $\operatorname{PG}(k-1,q)$, where $k\geq 3$, we have $n=r[k-1]_{q}$ for the cardinality and $s=r[k-2]_{q}+q^{k-2}$, $t=r[k-2]_{q}$ for the two different hyperplane multiplicities. Our first condition is
$11\leq r\leq[k]_{q}-$ 1 (34)
since since there are only $[k]_{q}$ different hyperplanes in $\operatorname{PG}(k-1,q)$, $r=0$ yields the trivial multiset of points with cardinality 00, and $r=[k]_{q}$ yields the multiset of points $[k-1]_{q}\cdot\chi_{V}$, which is a complement if the trivial multiset of points. For the canonical multiset of points we also need the parameters $\mu$ and $g$ to state
$n^{\prime}=\frac{1}{g}\cdot\left(r[k-1]_{q}-\mu[k]_{q}\right)$ (35)
for the cardinality and
$s_{0}=\frac{1}{g}\cdot\left(r[k-2]_{q}+q^{k-2}-\mu[k-1]_{q}\right),t_{0}=\frac{1}{g}\cdot\left(r[k-2]_{q}-\mu[k-1]_{q}\right)$ (36)
for the hyperplane multiplicities. Clearly, we have the conditions
$n^{\prime},s_{0}\in\mathbb{N}_{\geq_{1}}\text{ and }t_{0}\in\mathbb{N}$ (37)
as well as
$g\text{ divides }q^{k-2}$ (38)
and
$n^{\prime}<[k]_{q}$ (39)
since there are only $[k]_{q}$ different points in $PG(k-1,q)$. Let us denote the number of hyperplanes with multiplicity $s_{0}$ by $a_{s}$ and the number of hyperplanes with multiplicity $t_{0}$ by $a_{t}$, so that $a_{s}+a_{t}=[k]_{q}$ and $s_{0}a_{s}-t_{0}a_{t}=n^{\prime}[k-1]_{q}$. Thus, we have
$a_{s}$ $=$ $\frac{g}{q^{k-2}}\cdot\left(n^{\prime}[k-1]_{q}-t_{0}[k]_{q}\right)$ (40)
$a_{t}$ $=$ $[k]_{q}-a_{s},$ (41)
so that we can require
$a_{s},a_{t}\in\mathbb{N}_{\geq 1}.$ (42)
The, so-called, standard equations for sets of points in $\operatorname{PG}(k-1,q)$ are completed by
(${s_{0}\choose 2}a_{s}+{t_{0}\choose 2}a_{t}={n^{\prime}\choose 2}[k-2]_{q}.$ (43)
So, in order to enumerate all possible candidates for the parameters of two-character sets $\mathcal{M}^{\prime}$ in $\operatorname{PG}(k-1,q)$ we loop over all choices of $11\leq r\leq[k]_{q}-$ 1, $0\leq\mu\leq r[k-2]_{q}/[k-1]_{q}$, and all divisors $g$ of $q^{k-2}$, compute $n^{\prime}$, $s_{0}$, $t_{0}$, $a_{s}$, $a_{t}$, and check the mentioned conditions. We remark that the weights of the corresponding two-weight codes $c^{\prime}$ are given by $w_{1}=n^{\prime}-s_{0}$ and $w_{2}=n^{\prime}-t_{0}$, so that $C^{\prime}$ is $q^{k-2}/g$-divisible and projective.
In the subsequent sections we list the obtained candidates and mention corresponding constructions from the literature. It turns out that for the binary case our necessary conditions are also sufficient for $k\leq 5$. For larger dimensions we have to exclude a few cases and also leave a few cases open for $k\in\{11,12\}$.
Before we start to state the numerical details we would like to mention two parametric series of constructions (assuming that the parameters $k$ and $q$ are clear from the context). For $11\leq i\leq k-$ 1 we may consider the characteristic function $\mathcal{S}_{i}:=\chi_{S}$ of an $i$-space $S$ with parameters $n^{\prime}=[i]_{q}$, $s_{0}=[i]_{q}$, and $t_{0}=[i-1]_{q}$, see [16, Construction SU], as well as the complement $\chi_{V}-\chi_{S}$ with $n^{\prime}=[k]_{q}-[i]_{q}$, $s_{0}=[k-1]_{q}-[i]-1_{q}$, and $t_{0}=[k-1]_{q}-[i]_{q}$. Whenever the dimension $k$ of the ambient space is even we can consider the characteristic function $\mathcal{P}_{i}$ of partial $(k/2)$-spreads consisting of $i$ pairwise disjoint $(k/2)$-spaces, where $10<i<[k]_{q}/[k/2]_{q}=q^{k/2}+$ 1, see [16, Construction SU2]. The parameters are given by $n^{\prime}=i[k/2]_{q}$, $s_{0}=i[k/2-1]_{q}+q^{k/2-1}$, and $t_{0}=i[k/2-1]_{q}$. Note that $\mathcal{P}_{1}$ has the same parameters as $\mathcal{S}_{k/2}$, $\mathcal{P}_{[k]_{q}-1}$ has the same parameters as the complement of $\mathcal{S}_{k/2}$, and the complement of $\mathcal{P}_{i}$ has the same parameters as $\mathcal{P}_{q^{k/2}+1-i}$. In general we can have lots of different non-isomorphic two-character sets with the same parameters, so that different parametric constructions may also cover the same parameters. For a third possible series $\mathcal{K}_{i}$ we currently need to assume $q=2$ and an even dimension $k$ of the ambient space. Moreover, no explicit construction is known. However, the parameters $n^{\prime}=i\cdot\left(q^{k/2}+1\right)$, $s_{0}=i\cdot\left(q^{k/2-1}+1\right)$, and $t_{0}=i\cdot\left(q^{k/2-1}+1\right)-q^{k/2-1}$ show a clear pattern. In principle we may choose $0<i<[k/2]_{q}$ but it will turn out that rather small and rather large values of $i$ have to be excluded. Note that the complementary parameters of $\mathcal{K}_{i}$ are given by $\mathcal{K}_{[k/2]_{q}-i}$. We will discuss explicit constructions attaining the parameters of $\mathcal{K}_{i}$ when considering explicit dimensions $k\equiv 0\pmod{2}$. Finally, we remark that for a two-character set $\mathcal{M}$ the corresponding set $\overline{\mathcal{M}}$ of hyperplanes with the maximum hyperplane multiplicity is also a two-character set, so that interchanging the roles of $\mathcal{M}$ and $\overline{\mathcal{M}}$ gives another general construction.
A.1. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(3,2)}$
In Table 5 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(3,2)$. We have tried to arrange those parameters in several series so that patterns become visible. While parameters of the complementary set of parameters can be easily computed we will mostly state them nevertheless. The first block consists of the parameters for the multisets $\mathcal{S}_{i}$ for $1\leq i\leq 3$ obtained from the subspace construction and their complements. The second block consists of the parameters for the multisets $\mathcal{P}_{i}$ for $2\leq i\leq 3$ obtained from the partial spread construction. Note that $\mathcal{P}_{1}$ has the same parameters as $\mathcal{S}_{2}$, $\mathcal{P}_{4}$ has the same parameters as the complement of $\mathcal{S}_{2}$, and the complement of $\mathcal{P}_{i}$ has the same parameters as $\mathcal{P}_{5-i}$. The third block has the parameters of $\mathcal{K}_{i}$ for $1\leq i\leq 2$.
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction
7 3 4 1 1 0 7 8 [16, Construction SU]
3 1 2 3 3 1 3 12 [16, Construction SU]
1 0 1 7 7 3 1 14 [16, Construction SU]
14 6 1 8 4 0 14 1 complement
12 4 2 12 6 4 12 3 complement
8 0 4 14 7 6 8 7 complement
6 2 2 6 4 2 6 9 [16, Construction SU2]
9 3 2 9 5 3 9 6 complement
10 4 2 5 3 1 10 5 [16, Construction RT2]
5 1 2 10 6 4 5 10 complement
Table 5. Feasible parameters for two-character sets in $\operatorname{PG}(3,2)$.
A.2. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(4,2)}$
In Table 6 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(4,2)$. Here we have just one block with attaining examples $\mathcal{S}_{1},\dots,\mathcal{S}_{4}$ and their complements.
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction
15 7 8 1 1 0 15 16 [16, Construction SU]
7 3 4 3 3 1 7 24 [16, Construction SU]
3 1 2 7 7 3 3 28 [16, Construction SU]
1 0 1 15 15 7 1 30 [16, Construction SU]
30 14 1 16 8 0 30 1 complement
28 12 2 24 12 8 28 3 complement
24 8 4 28 14 12 24 7 complement
16 0 8 30 15 14 16 15 complement
Table 6. Feasible parameters for two-character sets in $\operatorname{PG}(4,2)$.
A.3. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(5,2)}$
In Table 7 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(5,2)$. The first block consists of the parameters for the multisets $\mathcal{S}_{i}$ for $1\leq i\leq 5$ obtained from the subspace construction and their complements. The second block consists of the parameters for the multisets $\mathcal{P}_{i}$ for $2\leq i\leq 7$ obtained from the partial spread construction. The third block has the parameters of $\mathcal{K}_{i}$ for $2\leq i\leq 5$. The parameters for $\mathcal{K}_{1}$ as well as the complementary parameters for $\mathcal{K}_{6}$ cannot be attained:
Lemma A.1.
No $4$-divisible $[9,k]_{2}$-code exists.
  • Proof. Apply e.g. [31, Theorem 1]. $\square$
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction/non-existence
31 15 16 1 1 0 31 32 [16, Construction SU]
15 7 8 3 3 1 15 48 [16, Construction SU]
7 3 4 7 7 3 7 56 [16, Construction SU]
3 1 2 15 15 7 3 60 [16, Construction SU]
1 0 1 31 31 15 1 62 [16, Construction SU]
62 30 1 32 16 0 62 1 complement
60 28 2 48 24 16 60 3 complement
56 24 4 56 28 24 56 7 complement
48 16 8 60 30 28 48 15 complement
32 0 16 62 31 30 32 31 complement
14 6 4 14 10 6 14 49 [16, Construction SU2]
21 9 4 21 13 9 21 42 [16, Construction SU2]
28 12 4 28 16 12 28 35 [16, Construction SU2]
35 15 4 35 19 15 35 28 complement
42 18 4 42 22 18 42 21 complement
49 21 4 49 25 21 49 14 complement
54 26 4 9 5 1 54 9 Lemma A.1
45 21 4 18 10 6 45 18 [2]
36 16 4 27 15 11 36 27 [16, Construction RT2]
27 11 4 36 20 16 27 36 complement
18 6 4 45 25 21 18 45 complement
9 1 4 54 30 26 9 54 complement does not exist
Table 7. Feasible parameters for two-character sets in $\operatorname{PG}(5,2)$.
A.4. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(6,2)}$
In Table 8 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(6,2)$. Here we have just one block with attaining examples $\mathcal{S}_{1},\dots,\mathcal{S}_{6}$ and their complements.
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction
63 31 32 1 1 0 63 64 [16, Construction SU]
31 15 16 3 3 1 31 96 [16, Construction SU]
15 7 8 7 7 3 15 112 [16, Construction SU]
7 3 4 15 15 7 7 120 [16, Construction SU]
3 1 2 31 31 15 3 124 [16, Construction SU]
1 0 1 63 63 31 1 126 [16, Construction SU]
126 62 1 64 32 0 126 1 complement
124 60 2 96 48 32 124 3 complement
120 56 4 112 56 48 120 7 complement
112 48 8 120 60 56 112 15 complement
96 32 16 124 62 60 96 31 complement
64 0 32 126 63 62 64 63 complement
Table 8. Feasible parameters for two-character sets in $\operatorname{PG}(6,2)$.
A.5. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(7,2)}$
In Table 9 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(7,2)$ that are not covered by $\mathcal{P}_{i}$ for $2\leq i\leq 15$, $\mathcal{S}_{i}$ for $1\leq i\leq 7$, or their complements. There remains one block corresponding to the parameters of $\mathcal{K}_{i}$ for $1\leq i\leq 14$. The parameters for $\mathcal{K}_{1},\mathcal{K}_{2}$ as well as the complementary parameters for $\mathcal{K}_{14},\mathcal{K}_{13}$ cannot be attained:
Lemma A.2.
No $8$-divisible $[17,k]_{2}$-code exists. Each $8$-divisible $[34,k]_{2}$-code has column multiplicity at least two.
  • Proof. For the first statement apply e.g. [31, Theorem 1] and for the second see e.g. [33] or apply [29, Theorem 12]. $\square$
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction/non-existence
238 118 8 17 9 1 238 17 Lemma A.2
221 109 8 34 18 10 221 34 Lemma A.2
204 100 8 51 27 19 204 51 [16, Construction CY1], [32]
187 91 8 68 36 28 187 68 [16, Construction FE1], [25], [32]
170 82 8 85 45 37 170 85 [16, Construction CY1], [32]
153 73 8 102 54 46 153 102 [16, Construction CY2], [32]
136 64 8 119 63 55 136 119 [16, Construction RT2], [32]
119 55 8 136 72 64 119 136 complement
102 46 8 153 81 73 102 153 complement
85 37 8 170 90 82 85 170 complement
68 28 8 187 99 91 68 187 complement
51 19 8 204 108 100 51 204 complement
34 10 8 221 117 109 34 221 complement does not exist
17 1 8 238 126 118 17 238 complement does not exist
Table 9. Feasible parameters for two-character sets in $\operatorname{PG}(7,2)$ excluding those for $\mathcal{P}_{i}$, $\mathcal{S}_{i}$, and their complements.
A.6. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(8,2)}$
In Table 10 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(8,2)$ that are not covered by $\mathcal{S}_{i}$ for $1\leq i\leq 8$, or their complements.
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction
441 217 8 196 100 84 441 70 [2]
73 33 8 219 123 107 73 438 [32]
438 214 8 292 148 132 438 73 complement
70 30 8 315 171 155 70 441 complement
315 155 16 70 38 30 315 196 [2]
219 107 16 73 41 33 219 292 [23], [32]
292 132 16 438 222 214 292 219 complement
196 84 16 441 225 217 196 315 complement
Table 10. Feasible parameters for two-character sets in $\operatorname{PG}(8,2)$ excluding those for $\mathcal{S}_{i}$ and their complements.
A.7. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(9,2)}$
In Table 11 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(9,2)$ that are not covered by $\mathcal{P}_{i}$ for $2\leq i\leq 31$, $\mathcal{S}_{i}$ for $1\leq i\leq 9$, or their complements. There remains one block corresponding to the parameters of $\mathcal{K}_{i}$ for $1\leq i\leq 30$. The parameters for $\mathcal{K}_{1},\dots,\mathcal{K}_{5}$ as well as the complementary parameters for $\mathcal{K}_{30},\dots,\mathcal{K}_{26}$ cannot be attained:
Lemma A.3.
No projective $16$-divisible $[n,k]_{2}$-code exists for $n\in\{33,66,99,132\}$.
  • Proof. For $n\in\{33,66,99\}$ we apply [29, Theorem 12]. For $n=132$ we apply [34, Corollary 6.15] with $t=4$. $\square$
Lemma A.4.
No $[165,10,80]_{2}$-code exists.
  • Proof. See [8, Corollary 1]. $\square$
The result is concluded from the non-existence of a $[69,9,32]_{2}$-code applied to the residual code of a codeword of weight $96$ c.f. [26, Section 3.4].
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction/non-existence
990 494 16 33 17 1 990 33 Lemma A.3
957 477 16 66 34 18 957 66 Lemma A.3
924 460 16 99 51 35 924 99 Lemma A.3
891 443 16 132 68 52 891 132 Lemma A.3
858 426 16 165 85 69 858 165 Lemma A.4
825 409 16 198 102 86 825 198 [32]
792 392 16 231 119 103 792 231 [20], [32]
759 375 16 264 136 120 759 264 [16, Construction FE1], [32]
726 358 16 297 153 137 726 297 [20], [32]
693 341 16 330 170 154 693 330 [20], [32]
660 324 16 363 187 171 660 363 [20], [32]
627 307 16 396 204 188 627 396 [20], [32]
594 290 16 429 221 205 594 429 [20], [32]
561 273 16 462 238 222 561 462 [20], [32]
528 256 16 495 255 239 528 495 [16, Construction RT2], [32]
495 239 16 528 272 256 495 528 complement
462 222 16 561 289 273 462 561 complement
429 205 16 594 306 290 429 594 complement
396 188 16 627 323 307 396 627 complement
363 171 16 660 340 324 363 660 complement
330 154 16 693 357 341 330 693 complement
297 137 16 726 374 358 297 726 complement
264 120 16 759 391 375 264 759 complement
231 103 16 792 408 392 231 792 complement
198 86 16 825 425 409 198 825 complement
165 69 16 858 442 426 165 858 complement does not exist
132 52 16 891 459 443 132 891 complement does not exist
99 35 16 924 476 460 99 924 complement does not exist
66 18 16 957 493 477 66 957 complement does not exist
33 1 16 990 510 494 33 990 complement does not exist
Table 11. Feasible parameters for two-character sets in $\operatorname{PG}(9,2)$ excluding those for $\mathcal{P}_{i}$, $\mathcal{S}_{i}$, and their complements.
A.8. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(10,2)}$
In Table 12 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(10,2)$ that are not covered by $\mathcal{S}_{i}$ for $1\leq i\leq 10$, or their complements.
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction/non-existence
1780 884 16 712 360 328 1780 267 open
276 132 16 759 407 375 276 1771 [16, Construction RT5]
1771 875 16 1288 648 616 1771 276 complement
267 123 16 1335 695 663 267 1780 complement of open case
1335 663 32 267 139 123 1335 712 open
759 375 32 276 148 132 759 1288 [16, Construction RT5]
1288 616 32 1771 891 875 1288 759 complement
712 328 32 1780 900 884 712 1335 complement of open case
Table 12. Feasible parameters for two-character sets in $\operatorname{PG}(10,2)$ excluding those for $\mathcal{S}_{i}$ and their complements.
We remark that a projective $32$-divisible binary code of length $712$ and a projective $16$-divisible binary code of length $267$ indeed exist, see [34]. However, the currently known constructions use more than two weights.
A.9. Feasible parameters of two-character sets in $\mathbf{\operatorname{PG}(11,2)}$
In Table 13 and Table 14 we have listed the feasible parameters of two-character sets in $\operatorname{PG}(9,2)$ that are not covered by $\mathcal{P}_{i}$ for $2\leq i\leq 63$, $\mathcal{S}_{i}$ for $1\leq i\leq 11$, or their complements. There remains one block corresponding to the parameters of $\mathcal{K}_{i}$ for $1\leq i\leq 62$. The parameters for $\mathcal{K}_{1},\dots,\mathcal{K}_{6}$ as well as the complementary parameters for $\mathcal{K}_{62},\dots,\mathcal{K}_{57}$ cannot be attained:
Lemma A.5.
No projective $32$-divisible $[n,k]_{2}$-code exists for $n\in\{65,130,195,260,325,390\}$.
  • Proof. For $n\in\{65,130,195,260\}$ we apply [29, Theorem 12]. For $n\in\{325,390\}$ we apply [34, Corollary 6.15] with $t=5,6$. $\square$
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction/non-existence
4030 2014 32 65 33 1 4030 65 Lemma A.5
3965 1981 32 130 66 34 3965 130 Lemma A.5
3900 1948 32 195 99 67 3900 195 Lemma A.5
3835 1915 32 260 132 100 3835 260 Lemma A.5
3770 1882 32 325 165 133 3770 325 Lemma A.5
3705 1849 32 390 198 166 3705 390 Lemma A.5
3640 1816 32 455 231 199 3640 455 [16, Construction CY1], [32]
3575 1783 32 520 264 232 3575 520 open
3510 1750 32 585 297 265 3510 585 open
3445 1717 32 650 330 298 3445 650 open
3380 1684 32 715 363 331 3380 715 open
3315 1651 32 780 396 364 3315 780 [2], [32]
- 3250 1618 32 845 429 397 3250 845 [32]
3120 1552 32 975 495 463 3120 975 [32]
3055 1519 32 1040 528 496 3055 1040 [32]
2990 1486 32 1105 561 529 2990 1105 [32]
2925 1453 32 1170 594 562 2925 1170 [32]
2860 1420 32 1235 627 595 2860 1235 open
2795 1387 32 1300 660 628 2795 1300 [32]
2730 1354 32 1365 693 661 2730 1365 [16, Construction CY1,CY2], [32]
2665 1321 32 1430 726 694 2665 1430 [32]
2600 1288 32 1495 759 727 2600 1495 [32]
2535 1255 32 1560 792 760 2535 1560 [32]
2470 1222 32 1625 825 793 2470 1625 [32]
2405 1189 32 1690 858 826 2405 1690 [32]
2340 1156 32 1755 891 859 2340 1755 [32]
2275 1123 32 1820 924 892 2275 1820 [16, Construction CY1,CY2], [32]
2210 1090 32 1885 957 925 2210 1885 [32]
2145 1057 32 1950 990 958 2145 1950 [32]
2080 1024 32 2015 1023 991 2080 2015 [16, Construction RT2], [32]
2015 991 32 2080 1056 1024 2015 2080 complement
1950 958 32 2145 1089 1057 1950 2145 complement
1885 925 32 2210 1122 1090 1885 2210 complement
1820 892 32 2275 1155 1123 1820 2275 complement
1755 859 32 2340 1188 1156 1755 2340 complement
1690 826 32 2405 1221 1189 1690 2405 complement
1625 793 32 2470 1254 1222 1625 2470 complement
1560 760 32 2535 1287 1255 1560 2535 complement
1495 727 32 2600 1320 1288 1495 2600 complement
1430 694 32 2665 1353 1321 1430 2665 complement
1365 661 32 2730 1386 1354 1365 2730 complement
1300 628 32 2795 1419 1387 1300 2795 complement
1235 595 32 2860 1452 1420 1235 2860 complement of open case
1170 562 32 2925 1485 1453 1170 2925 complement
1105 529 32 2990 1518 1486 1105 2990 complement
1040 496 32 3055 1551 1519 1040 3055 complement
975 463 32 3120 1584 1552 975 3120 complement
845 397 32 3250 1650 1618 845 3250 complement
780 364 32 3315 1683 1651 780 3315 complement
Table 13. Feasible parameters for two-character sets in $\operatorname{PG}(11,2)$ excluding those for $\mathcal{P}_{i}$, $\mathcal{S}_{i}$, and their complements – part 1.
$r$ $\mu$ $g$ $n^{\prime}$ $s_{0}$ $t_{0}$ $a_{s}$ $a_{t}$ construction/non-existence
715 331 32 3380 1716 1684 715 3380 complement of open case
650 298 32 3445 1749 1717 650 3445 complement of open case
585 265 32 3510 1782 1750 585 3510 complement of open case
520 232 32 3575 1815 1783 520 3575 complement of open case
455 199 32 3640 1848 1816 455 3640 complement
390 166 32 3705 1881 1849 390 3705 complement does not exist
325 133 32 3770 1914 1882 325 3770 complement does not exist
260 100 32 3835 1947 1915 260 3835 complement does not exist
195 67 32 3900 1980 1948 195 3900 complement does not exist
130 34 32 3965 2013 1981 130 3965 complement does not exist
65 1 32 4030 2046 2014 65 4030 complement does not exist
2808 1400 64 234 122 106 2808 1287 [16, Construction FE3], [17], [32]
1400 696 64 245 133 117 1400 2695 open
2295 1143 64 270 142 126 2295 1800 [2], [32]
1911 951 64 273 145 129 1911 2184 [17], [32]
2184 1032 64 3822 1918 1902 2184 1911 complement
1800 840 64 3825 1921 1905 1800 2295 complement
2695 1287 64 3850 1930 1914 2695 1400 complement of open case
1287 583 64 3861 1941 1925 1287 2808 complement
3185 1585 32 910 462 430 3185 910 [16, Construction CY1,CY2]
910 430 32 3185 1617 1585 910 3185 complement
[…]
245 117 16 1400 760 696 245 3850 open
3825 1905 16 1800 904 840 3825 270 [2], [32]
273 129 16 1911 1015 951 273 3822 [18], [32]
3822 1902 16 2184 1096 1032 3822 273 complement
270 126 16 2295 1207 1143 270 3825 complement
3850 1914 16 2695 1351 1287 3850 245 complement of open case
234 106 16 2808 1464 1400 234 3861 complement
Table 14. Feasible parameters for two-character sets in $\operatorname{PG}(11,2)$ excluding those for $\mathcal{P}_{i}$, $\mathcal{S}_{i}$, and their complements – part 2.
W.r.t. the series $\mathcal{K}_{i}$ we remark that it is currently unknown whether projective $32$-divisible $[n,k]_{2}$-codes exist for $n\in\{520,585,650,715\}$, see e.g. [34]. W.r.t. the other open cases no construction for a projective $\Delta$-divisible $[n,k]_{2}$-code is known for $(\Delta,n)=(16,245)$ and $(\Delta,n)=(64,1400)$.